### v1

# Hilbert’s Theorem 90, periodicity, and roots of Artin-Schreier polynomials

Abstract
Let $E/F$ be a cyclic field extension of degree $n$, and let $\sigma$ generate the group $\textup{Gal}(E/F)$. If $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}\sigma^{i}y=0$, then the additive form of Hilbert’s Theorem 90 asserts that $y=\sigma x-x$ for some $x\in E$. Suppose that $E$ has characteristic $p$. We prove that $x$ gives rise to a periodic sequence $x_{0},x_{1},\dots$ which has period $pn_{p}$, where $n_{p}$ is the largest $p$-power that divides $n$. As an application, we find closed-form expressions for the roots of Artin-Schreier polynomials $t^{p}-t-y$. Let $y$ lie in the finite field ${\mathbb{F}}_{p^{n}}$ of order $p^{n}$. The Artin-Schreier polynomial $t^{p}-t-y\in{\mathbb{F}}_{p^{n}}[t]$ is reducible precisely when $\sum_{i=0}^{n-1}y^{p^{i}}=0$. In this case, $t^{p}-t-y=\prod_{k=0}^{p-1}(t-x-k)$ where $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ for some $z\in{\mathbb{F}}_{p^{e}}$ and $e=n_{p}$. The sequence $\left(\sum_{j=0}^{i-1}z^{p^{j}}\right)_{i\geqslant 0}$ is periodic with period $pe$, and if $e$ is small, then we give explicit $z$. Keywords: roots, reducible, Artin-Schreier polynomial, linearized polynomials 2020 Mathematics Subject Classification: 11T06, 11T55, 12E20
1 Introduction
This note began with the aims to show that there are infinitely many ‘quadratic formulas’ for quadratics over the finite field ${\mathbb{F}}_{2^{n}}$. The formulas have a periodicity, and the periodicity is related to the additive form of Hilbert’s Theorem 90 as we now explain.
Let $E/F$ be a cyclic field extension of degree $n$, and let $\sigma$ generate the Galois group $\textup{Gal}(E/F)$. If $\sum_{i=0}^{n-1}\sigma^{i}y=0$, then the additive form of Hilbert’s Theorem 90 asserts that $y=\sigma x-x$ for some $x\in E$. In this note, $E$ has prime characteristic $p$, and $n_{p}$ denotes the largest power of $p$ that divides $n$. We prove that $x$ can be chosen to have the form $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\sigma^{j}z\sigma^{i}y$ where $z$ is fixed by $\sigma^{n_{p}}$, and satisfies $\sum_{i=0}^{n-1}\sigma^{i}z=1$. Furthermore, $x=\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ where $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ defines a periodic sequence $x_{0},x_{1},\dots$ with period $pn_{p}$, see Theorem 3.1(b,c). A standard proof of the additive form of Hilbert’s Theorem 90 e.g. [4, p. 215] uses the ‘standard’ expression $x^{\prime}=R(z,y)$, see Lemma 3.2, where $y=x^{\prime}-\sigma(x^{\prime})$. This standard expression is very natural in one sense, see the preamble to Lemma 3.2. However, it makes questions about periodicity unnatural. Lemma 3.2 relates $x=R(y,z)$ to $x^{\prime}=R(z,y)$ via the equation $x+x^{\prime}=-\sum_{i=0}^{n-1}\sigma^{i}(yz)$.
In some cases the generator $\sigma\colon E\to E$ may be taken to be the $p$th power map $\sigma(x)=x^{p}$, e.g. if $E={\mathbb{F}}_{p^{n}}$ is the finite field of order $p^{n}$ and $F={\mathbb{F}}_{p}$. If $\sum_{i=0}^{n-1}y^{p^{i}}=0$ and $\sum_{i=0}^{n-1}z^{p^{i}}=1$, then $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ is a root of the Artin-Schreier polynomial $t^{p}-t-y$. For example, if $n_{p}=1$, we may take $z=\frac{1}{n}$ and $x_{i}=\frac{i}{n}$ defines a sequence of period $p$. Furthermore, $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{p^{i}}$ is a closed-form expression for a root of $t^{p}-t-y$ in $E$ when $\sum_{i=0}^{n-1}y^{p^{i}}=0$. The case $n_{p}=p$ uses a different choice for $z$, and $n_{p}=p^{2}$ a different choice, etc. This suggests that there are infinitely many formulas for the roots of the polynomial $t^{p}-t-y$ over the finite field ${\mathbb{F}}_{p^{n}}$ that depend on the $p$-part $n_{p}$ of $n$. Choosing $z$ in the field ${\mathbb{F}}_{p^{e}}$ with $e=n_{p}$ and $\sum_{i=0}^{n-1}z^{p^{i}}=1$ is easy in practise: with probability $1-\frac{1}{p}$ a random element $\zeta\in{\mathbb{F}}_{p^{e}}$ has $\tau\coloneq\sum_{i=0}^{n-1}\zeta^{p^{i}}\neq 0$, so take $z=\tau^{-1}\zeta$. If $n_{p}$ is fixed, and we know $z\in{\mathbb{F}}_{p^{e}}$ satisfying $\sum_{i=0}^{n-1}z^{p^{i}}=1$, then $R(y,z)$ may be viewed as a closed-form expression for a root of $t^{p}-t-y$. For examples with $p=2$ see Lemma 4.8.
2 Background
This work fits into a very classical setting, namely generalizing Galois theory to characteristic $p$. The following important result follows from the additive form of Hilbert’s Theorem 90, see [2, Theorem 4.33] and [4, VIII, Theorem 11].
Theorem 2.1 (Artin-Schreier).
Let $F$ be field of prime characteristic $p$.
  1. (a)
    Let $E/F$ be a cyclic extension of degree $p$. Then $E=F(x)$ where $x^{p}-x\in F\setminus\{0\}$.
  2. (b)
    Given $y\in F$, the polynomial $t^{p}-t-y$ is either irreducible over $F$, or it has $p$ distinct roots $x+a$ where $a\in{\mathbb{F}}_{p}$ and $x\in F$.
The trace map $\textup{Tr}_{F}^{E}(x)=\sum_{\alpha\in\textup{Gal}(E/F)}\alpha(x)$ of a finite Galois extension $E/F$, is an $F$-linear map $E\to F$ which is nonzero by Artin’s Lemma [4, VIII, Theorem 7]. In certain cases, the trace map can be used to distinguish whether or not $t^{p}-t+y$ is irreducible.
Lemma 2.2.
Let $E/F$ be a cyclic Galois extension where $|E:F|=\textup{char}(F)=p$. Suppose that $\textup{Gal}(E/F)=\langle\sigma\rangle$ where $\sigma\colon E\to E$ is the $p$th power map $\sigma(x)=x^{p}$. For $y\in E$, the polynomial $t^{p}-t+y$ is reducible over $E$ if and only if $\textup{Tr}_{F}^{E}(y)=\sum_{i=0}^{p-1}y^{p^{i}}=0$. Furthermore, $E=F(z)$ for some $z\in E$ where $z^{p}-z^{p-1}+1=0$ and $\textup{Tr}^{E}_{F}(z)=1$.
Proof.
If $\textup{Tr}_{F}^{E}(y)=0$, then by Hilbert’s Theorem 90, $y=z-\sigma(z)=z-z^{p}$ for some $z\in E$. Hence $t^{p}-t+y$ is reducible over $E$. Conversely, suppose if $t^{p}-t+y$ is reducible over $E$, then $t^{p}-t+y$ has a root $z\in E$, by Theorem 2.1(b). Using $\sigma z=z^{p}$ shows
$0=\textup{Tr}^{E}_{F}(0)=\textup{Tr}^{E}_{F}(z^{p}-z+y)=\textup{Tr}^{E}_{F}(\sigma z)-\textup{Tr}^{E}_{F}(z)+\textup{Tr}^{E}_{F}(y)=\textup{Tr}^{E}_{F}(y).$
Finally, set $y=1$. Then $\textup{Tr}_{F}^{E}(1)=p\cdot 1=0$ and some $\zeta\in E$ satisfies $t^{p}-t+1=0$. However, $\sigma\zeta=\zeta^{p}=\zeta-1$ implies $\sigma^{i}\zeta=\zeta^{p}=\zeta-i$ so $\zeta$ has $p$ distinct conjugates in $E$ and so $E=F(\zeta)$. Therefore $E=F(z)$ where $z=\zeta^{-1}$ is a root of the (reverse) polynomial $t^{p}-t^{p-1}+1$. Since the conjugates $z,\sigma z,\dots,\sigma^{p-1}z$ are pairwise distinct, $t^{p}-t^{p-1}+1$ is also irreducible over $F$. The coefficient of $t^{p-1}$ shows $\textup{Tr}^{E}_{F}(z)=1$, as claimed. ∎
[…]
If $E={\mathbb{F}}_{2^{n}}$ is a finite field of order $2^{n}$, then $x^{2}+bx+c$ has a double root $x=c^{2^{n-1}}$ if $b=0$ because $c^{2^{n}}=c$ holds. If $b\neq 0$, then $x^{2}+bx+c=b^{-2}\left(t^{2}+t+b^{-2}c\right)$ where $t=b^{-1}x$. Taking $y=b^{-2}c$, the quadratic $t^{2}+t+y$ has roots $R(y,z),R(y,z)+1$ where $R(y,z)$ is given by the periodic formula $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{2^{j}}y^{2^{i}}$, see Lemma 4.8.
3 Periodicity and Hilbert’s additive Theorem 90
If $E/F$ is a finite Galois extension and $\sigma\in\textup{Gal}(E/F)$, then $\textup{Tr}^{E}_{F}(\sigma x)=\textup{Tr}^{E}_{F}(x)$ holds. Hence if $y\in E$ has the form $\sigma x-x$ for some $x\in E$ and $\sigma\in\textup{Gal}(E/F)$, then $\textup{Tr}^{E}_{F}(y)=0$ holds. If $\textup{Gal}(E/F)$ is finite and cyclic, then the converse holds by Hilbert’s Theorem 90. Throughout this section $E/F$ is a cyclic Galois extension of degree $n$ and prime characteristic $p$, and $\textup{Gal}(E/F)$ is generated by $\sigma$. Hence $\sigma^{n}=1$ and $\textup{Tr}_{F}^{E}(y)=\sum_{i=0}^{n-1}\sigma^{i}y$.
We prove in Theorem 3.1 that the element $x$ has an intrinsic periodicity. This observation is used in Corollary 4.1 to give a polynomial formula for the roots of (reducible) Artin-Schreier polynomials.
Theorem 3.1.
Let $E/F$ be a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\textup{char}(F)=p$. Let $n$ have $p$-part $n_{p}$, and $p^{\prime}$-part $n_{p^{\prime}}=n/n_{p}$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=0$, then
  1. (a)
    $y=\sigma x-x$ where $x=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$ for some $z\in E$ with $\textup{Tr}^{E}_{F}(z)=1$;
  2. (b)
    $F(z)$ contains the fixed field of $\sigma^{n_{p}}$, and $z$ may be chosen so that $|F(z):F|=n_{p}$;
  3. (c)
    the sequence $(\sum_{j=0}^{i-1}\sigma^{j}z)_{i\geqslant 0}$ is periodic with period $pe$ where $e=|F(z):F|$ and $n_{p}\mid e$.
Proof.
(a) Suppose that $x$ has the form $\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ where $x_{0},\dots,x_{n-1}\in E$. Then $\sigma x-x=-x_{0}y+\sum_{i=1}^{n-1}(\sigma x_{i-1}-x_{i})\sigma^{i}y+\sigma x_{n-1}\sigma^{n}y$. Assume for the moment that there exists an element $z\in E$ satisfying
(1) $\sigma x_{i-1}-x_{i}=-z\quad\textup{for $1\leqslant i<n$, and}\quad\sigma x_{n-1}-x_{0}=1-z.$
Substituting these equations and using $\sigma^{n}y=y$ gives
$\sigma x-x=-x_{0}y-z\sum_{i=1}^{n-1}\sigma^{i}y+(1-z+x_{0})y=y-z\sum_{i=0}^{n-1}\sigma^{i}y=y$
as desired. Next, assume that $z$ and $x_{0}$ are given. The equations $\sigma x_{i-1}-x_{i}=-z$ show that $x_{i}=\sigma^{i}x_{0}+\sigma^{i-1}z+\cdots+\sigma z+z$ holds for $1\leqslant i<n$. Rewrite the $n$ equations (1) as $\sigma^{n-i+1}x_{i-1}-\sigma^{n-i}x_{i}=-\sigma^{n-i}z$ for $1\leqslant i<n$ and $\sigma x_{n-1}-x_{0}=1-z$. Adding these equations gives $0=\sigma^{n}x_{0}-x_{0}=1-\textup{Tr}_{F}^{E}(z)$. Hence $\textup{Tr}_{F}^{E}(z)=1$ must hold.
We prove that these necessary equations are sufficient. First, the trace map $\textup{Tr}_{F}^{E}\colon E\to F$ is surjective, so we may choose $z\in E$ so that $\textup{Tr}_{F}^{E}(z)=1$. Now define $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where the empty sum $x_{0}=\sum_{j=1}^{0}\sigma^{j}z$ is 00 by convention. Then the $n$ equations in (1) hold, and $\sigma x-x$ equals $y$.
(b) First, we show that the fixed subfield $K$ of $\sigma^{n_{p}}$ contains an element $z$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Since $n_{p^{\prime}}\neq 0$ in $F$ and the map $\textup{Tr}^{K}_{F}\colon K\to F$ is surjective, there exists an element $z\in K$ such that $\textup{Tr}^{K}_{F}(z)=(n_{p^{\prime}})^{-1}$. Then $\textup{Tr}^{E}_{K}(z)=|E:K|z=n_{p^{\prime}}z$ and hence $\textup{Tr}^{E}_{F}(z)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(z))=n_{p^{\prime}}\textup{Tr}^{K}_{F}(z)=1$. Next, suppose that $L=F(z)$ and $\textup{Tr}^{E}_{F}(z)=1$. Then $1=\textup{Tr}^{L}_{F}(\textup{Tr}^{E}_{L}(z))=|E:L|\textup{Tr}^{L}_{F}(z)$. Therefore $p\nmid|E:L|$, so $n_{p}=|K:F|$ divides $e=|L:F|$ as claimed. In summary, the minimal polynomial over $F$ of an element $z$ with $\textup{Tr}^{E}_{F}(z)=1$ has degree a multiple of $n_{p}$, and degree precisely $n_{p}$ if $z$ lies in $K\setminus K_{0}$ where $K_{0}$ is the unique (maximal) subfield of $K$ with $F\leqslant K_{0}<K$ and $|K:K_{0}|=p$.
(c) Let $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where $\textup{Tr}^{E}_{F}(z)=1$. Let $L=F(z)$ and $e=|L:F|$. Hence $x_{e}=\textup{Tr}^{L}_{F}(z)\neq 0$ by part (b) and $\sigma^{j}z=z$ implies $e\mid j$. Suppose that $x_{k}=0$ where $0\leqslant k<pe$ and write $k=ie+j$ where $0\leqslant i<p$ and $0\leqslant j<e$. The equation $x_{ie+j}=ix_{e}+x_{j}$ shows that $x_{0}=x_{pe}=0$. Thus $x_{k}=ix_{e}+x_{j}=0$ and $x_{e}\in F$ gives
$x_{j+1}-z=\sigma x_{j}=\sigma(-ix_{e})=-ix_{e}=x_{j}.$
Whence $\sigma^{j}z=x_{j+1}-x_{j}=z$. This implies that $e\mid j$ and hence $j=0$. However, $0=x_{k}=x_{ie}=ix_{e}$ shows that $0=i=j=k$. Therefore $x_{\ell}\neq 0$ holds for $0<\ell<pe$, and the sequence $x_{0},x_{1},\dots$ is periodic, with period precisely $pe$. ∎
The attentive reader will notice that the proof of Theorem 3.1(a) differs from the standard proof: a symmetry has been broken as we now explain. There are two versions of theorems regarding the kernels of the norm and trace maps. The former is usually called Hilbert’s (multiplicative) Theorem 90 because of the numbering in [1, p. 272], and the latter the additive form of Hilbert’s Theorem 90. Emmy Noether generalized the method used to prove these theorems, essentially by introducing 1-cocyles, to handle a noncyclic (finite) Galois group $G\coloneq\textup{Gal}(E/F)$, see [2, Theorems 4.29 and 4.30]. Her first result [2, Theorem 4.29] says: given a function $y\colon G\to E\setminus\{0\}$ satisfying $y_{\sigma\tau}=y_{\sigma}\sigma(y_{\tau})$ for $\sigma,\tau\in G$, there exists an $x\in E\setminus\{0\}$ such that $y_{\tau}=x\tau(x)^{-1}$. Her second result [2, Theorem 4.30] says: given a function $y\colon G\to E$ satisfying $y_{\sigma\tau}=\sigma(y_{\tau})+y_{\sigma}$ for $\sigma,\tau\in G$, there exists an $x\in E$ such that $y_{\tau}=x-\tau(x)$. When $G=\langle\sigma\rangle$ is cyclic order $n$, the 1-cocycle conditions suggest the definitions $y_{\phi^{i}}=y(\sigma y)\cdots(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$, and similarly in the additive case $y_{\phi^{i}}=y+(\sigma y)+\cdots+(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{0}=y$. This gives rise to the (multiplicative and additive) formulas $x=\sum_{i=0}^{n-1}\left(\prod_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$ and $x=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$, respectively. The additive expression is $R(z,y)$ where $R(y,z)$ is in Theorem 3.1(a). The expressions $R(z,y)$ and $R(y,z)$ are related as follows.
Lemma 3.2.
Let $E/F$ be a cyclic Galois extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$. If $y,z\in E$ have traces $\textup{Tr}^{E}_{F}(y)=0$ and $\textup{Tr}^{E}_{F}(z)=1$, then $R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$ satisfies $R(y,z)=-\textup{Tr}^{E}_{F}(yz)-R(z,y)$, $\sigma R(y,z)-R(y,z)=y$ and $R(z,y)-\sigma R(z,y)=y$.
Proof.
The equations $\sum_{j=0}^{n-1}\sigma^{j}z=1$ and $\sum_{i=0}^{n-1}\sigma^{i}y=0$ imply
$R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y=\sum_{i=0}^{n-1}\left(1-\sum_{j=i}^{n-1}\sigma^{j}z\right)\sigma^{i}y=-\sum_{i=0}^{n-1}\left(\sum_{j=i}^{n-1}\sigma^{j}z\right)\sigma^{i}y.$
Swapping the order of summation, swapping $i\leftrightarrow j$, and noting $\textup{Tr}^{E}_{F}(zy)=\sum_{i=0}^{n-1}\sigma(zy)$ gives
$R(y,z)=-\sum_{j=0}^{n-1}\sum_{i=0}^{j}\sigma^{j}z\sigma^{i}y=-\sum_{i=0}^{n-1}\sum_{j=0}^{i}\sigma^{i}z\sigma^{j}y=-\textup{Tr}^{E}_{F}(zy)-\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z.$
Therefore $R(y,z)=-\textup{Tr}^{E}_{F}(yz)-R(z,y)$. Theorem 3.1 proves that $\sigma R(y,z)-R(y,z)=y$ holds. Since $\sigma\textup{Tr}^{E}_{F}(yz)=\textup{Tr}^{E}_{F}(yz)$ it follows that $R(z,y)-\sigma R(z,y)=y$, as claimed. ∎
Remark 3.3.
Surprisingly, Hilbert did not prove the additive form of Hilbert’s Theorem 90. Lang in [4, p. 215] gives a proof, which he attributes to Emil Artin. It uses the expression $R(z,y)$. By contrast Theorem 3.1(a) uses $R(y,z)$, and the form $R(y,z)=\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ with $x_{i}=\sum_{j=0}^{n-1}\sigma^{j}z$ suggests the periodicity result in parts (b,c). The next section considers closed-form expressions for the roots of Artin-Schreier polynomials.
4 Formulas for the roots of Artin-Schreier polynomials
Let $E$ be a field of prime characteristic $p$. The $p$th power map $x\mapsto x^{p}$ defines a monomorphism $E\to E$. In this section, we assume that ${\mathbb{F}}_{q}\leqslant F\leqslant E$ where $q=p^{f}$ and $E/F$ is cyclic of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$. One such example involves finite fields $E={\mathbb{F}}_{q^{n}}$ and $F={\mathbb{F}}_{q}$, another involves function fields $E={\mathbb{F}}_{q^{n}}(T)$ and $F={\mathbb{F}}_{q}(T)$.
Corollary 4.1.
Suppose $E/F$ is a cyclic Galois extension of degree $n$ where $\textup{char}(F)=p$ and $\sigma(x)=x^{q}$, $q=p^{f}$, generates $\textup{Gal}(E/F)$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}y^{q^{i}}=0$, then
  1. (a)
    $x=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}z^{q^{j}}\right)y^{q^{i}}$ is a root of $t^{q}-t-y$ where $z\in E$ satisfies $\textup{Tr}^{E}_{F}(z)=1$;
  2. (b)
    $F(z)$ contains the fixed field of $\sigma^{n_{p}}$, and $z$ may be chosen so that $|F(z):F|=n_{p}$;
  3. (c)
    the sequence $(\sum_{j=0}^{i-1}z^{q^{j}})_{i\geqslant 0}$ is periodic with period $pe$ where $e=|F(z):F|$.
In the case when $n$ is coprime to $p$, i.e. $n_{p^{\prime}}=1$, the formula above for $x$ simplifies.
Corollary 4.2.
Consider Corollary 4.1 when $p\nmid n$, that is $n_{p}=1$. If $y\in E$ has trace $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}y^{q^{i}}=0$, then $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$ satisfies $x^{q}-x-y=0$.
Proof.
By Corollary 4.1, we may choose $z$ to lie in $F$ and so $1=\textup{Tr}^{E}_{F}(z)=nz$. Hence $z=\frac{1}{n}$ and $x=R(y,z)=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$ satisfies $x^{q}-x-y=0$ by Theorem 4.1(a). ∎
Remark 4.3.
Corollary 4.2 shows that the roots of $t^{q}-t-y$ lie in the ${\mathbb{F}}_{q}$-subspace $\langle y,y^{q},\dots,y^{q^{n-1}}\rangle$ of ${\mathbb{F}}_{q^{n}}$ when $\textup{char}(F)=p\nmid n$. Since $\sum_{i=0}^{n-1}y^{q^{i}}=0$, the roots of $t^{q}-t-y$ lie in the (potentially larger) ${\mathbb{F}}_{q}$-subspace $\{x\in E\mid\textup{Tr}^{E}_{F}(x)=0\}$ of $E$ of codimension 1.
In the case when $n_{p}=p$, i.e. $p\mid n$ and $p^{2}\nmid n$, the element $z$ in Corollary 4.1 can be chosen to lie in the fixed field of $\sigma^{p}$, and to have a specific minimal polynomial.
Corollary 4.4.
Let $E/F$ be a cyclic field extension of degree $n$ where $n_{p}=\textup{char}(F)=p$. Suppose that $\textup{Gal}(E/F)=\langle\sigma\rangle$ where $\sigma\colon E\to E$ is the $p$th power map $\sigma(x)=x^{p}$. If $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}y^{q^{i}}=0$, and $\zeta$ satisfies $\zeta^{p}-\zeta^{p-1}+1=0$. Then $x^{p}-x-y=0$ where $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$.
Proof.
It follows from $n_{p}=p$, that $n_{p^{\prime}}=n/p$ is coprime to $p$. Let $K$ be the fixed field of $\sigma^{p}$. Then $K$ contains an element $\zeta$ satisfying $\zeta^{p}-\zeta^{p-1}+1=0$ and $\textup{Tr}^{L}_{F}(\zeta)=1$ by Lemma 2.2. Therefore $\textup{Tr}^{E}_{F}(\zeta)=n/p$ and $\textup{Tr}^{E}_{F}(z)=1$ where $z=(n/p)^{-1}\zeta$. By Theorem 2.1(a), $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$ satisfies $x^{p}-x-y=0$. ∎
We shall now apply Corollary 4.1 in the case that $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$, so that $E/F$ is cyclic and $\textup{Gal}(E/F)=\langle\sigma\rangle$ has order $n$ where $\sigma(x)=x^{p}$ is the $p$th power map.
Let $p,r$ be distinct primes and set $q=p^{f}$. Denote by $\textup{ord}_{r}(q)$ the smallest positive integer $e$ such that $q^{e}\equiv 1\pmod{r}$ and note that $e\mid(r-1)$. If $r\geqslant 2$, then the cyclotomic polynomial $\Phi_{r}(t)$ of degree $\phi(r)$ factors over ${\mathbb{F}}_{p}$ into irreducible polynomials each of degree $\textup{ord}_{r}(p)$ by [5, Theorem 2.47].
Corollary 4.1 shows that $z$ may be chosen to lie in a subfield of order $p^{n_{p}}$. Hence $z$ is a $(p^{n_{p}}-1)$th root of unity. Lemma 4.5 proves that, in many cases, $z$ may be chosen to have prime order $r\neq p$. This choice of $z$ may increase the period of the sequence $x_{0},x_{1},\dots$.
Lemma 4.5.
Let $p,r$ be unequal primes where $\textup{ord}_{r}(p)\kern-1.0pt=\kern-1.0pte$. Let $E\kern-1.0pt=\kern-1.0pt{\mathbb{F}}_{p^{n}}$, $K\kern-1.0pt=\kern-1.0pt{\mathbb{F}}_{p^{e}}$, $F\kern-1.0pt=\kern-1.0pt{\mathbb{F}}_{p}$ where $e\mid n$. There exists a primitive $r$th root of unity $\zeta\in K$ for which $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. If $p\nmid\frac{n}{e}$, and $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=0$, then $x=(\frac{n\tau}{e})^{-1}\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\zeta^{p^{j}}\right)y^{p^{i}}$ satisfies $x^{p}-x-y=0$. Finally, the sequence $(\sum_{j=0}^{i-1}\zeta^{p^{j}})_{i\geqslant 0}$ is periodic with period $ep$ where $n_{p}\mid e$.
Proof.
First, $r$ divides $p^{d}-1$ if and only if $e\mid d$. As $e\mid n$, all primitive $r$th roots of unity lie in $K$, and $F\leqslant K\leqslant E$ holds. The cyclotomic polynomial $\Phi_{r}(t)=t^{r-1}+t^{r-2}+\cdots+t+1$ factors over $F$ as a product of $(r-1)/e$ irreducible polynomials each of degree $e$, so the coefficient of $t^{e-1}$ must be nonzero in at least one irreducible. Hence some $r$th root of unity $\zeta\in K$ has $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. Thus $\textup{Tr}^{E}_{K}(\zeta)=\frac{n}{e}\zeta$, so that $\textup{Tr}^{E}_{F}(\zeta)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(\zeta))=\frac{n}{e}\textup{Tr}^{K}_{F}(\zeta)=\frac{n\tau}{e}\neq 0$. Therefore $z=(\frac{n\tau}{e})^{-1}\zeta\in K^{\times}$ satisfies $\textup{Tr}^{E}_{F}(z)=1$. The formula for $x=R(y,z)$ and the periodicity claim follow from Theorem 3.1(a,c). ∎
Corollary 4.6.
Let $p$ be a prime where $p\equiv 2\pmod{3}$. Let $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$ where $n$ is even and let $\omega\in E$ satisfy $\omega^{2}+\omega+1=0$. If $p\nmid(n/2)$, and $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=0$, then the sum $R(y)=(\frac{n}{2})^{-1}\sum_{i=0}^{n-1}\left(\lfloor\frac{i}{2}\rfloor-r\omega\right)y^{p^{i}}$ satisfies $R(y)^{p}-R(y)-y=0$ where $r\in\{0,1\}$ and $i\equiv r\pmod{2}$. Finally, the sequence $x_{0},x_{1},\dots$ is periodic of period $2p$ where $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r\omega\right)$.
Proof.
The condition $p\equiv 2\pmod{3}$ is equivalent to $\textup{ord}_{3}(p)=2$. Thus we may use Lemma 4.5 with $r=3$ and $\zeta$ replaced with $\omega$. However, $\omega^{p^{k}}$ is $\omega$ if $k$ is even, and $\omega^{2}$ if $k$ is odd. If $i=0,1,2,3,4,5,6$, then $s_{i}=\sum_{k=0}^{i-1}\omega^{p^{k}}$ equals 00, $\omega$, $\omega+\omega^{2}=-1$, $-1+\omega$, $-2$, $-2+\omega$, $-3$, etc. Therefore $s_{i}=-(i-r)/2+r\omega$ where $r\equiv i\pmod{2}\in\{0,1\}$ and $r\in\{0,1\}$. The values of $x_{i}=(\frac{n}{2})^{-1}\left(\frac{i-r}{2}-r\omega\right)$ are periodic of period $2p$ by Theorem 3.1(c). ∎
Remark 4.7.
If $n_{p}$ is bounded, then Corollary 4.1 gives a closed-form expression $R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}z^{q^{j}}\right)y^{q^{i}}$ for a root of a reducible polynomial $t^{q}-t-y$ over ${\mathbb{F}}_{q^{n}}$. Further, the element $z$ may be chosen to have degree $n_{p}$, the $p$-part of $n$, and the sequence $\left(\sum_{j=0}^{i-1}z^{q^{j}}\right)_{i\geqslant 0}$ has period $pn_{p}$. If $n_{p}$ is unbounded, then $R(y,z)$ is not a closed-form expression as it involves infinitely many choices for an element $z$ satisfying $\textup{Tr}^{E}_{F}(z)=1$.
The following result gives infinitely many ‘quadratic formulas’ over the finite field ${\mathbb{F}}_{2^{n}}$.
Lemma 4.8.
Let $y\in{\mathbb{F}}_{2^{n}}$ where $\sum_{i=0}^{n-1}y^{2^{i}}=0$. Then $x=\sum_{i=0}^{n-1}x_{i}y^{2^{i}}$ is a root of $t^{2}+t+y$ where $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ and $z\in{\mathbb{F}}_{2^{n}}$ has degree $n_{2}$. Write $z=(n/n_{2})^{-1}\zeta$. Table 4.8 lists choices for $\zeta$ and the minimal polynomial $m_{\zeta/{\mathbb{F}}_{2}}(t)$ of $\zeta$, as a function of the value of $n_{2}$.
$n_{2}$ $1$ $2$ $4$ $8$ $16$cdots
$\zeta$ $1$ $\omega$ $\alpha$ $\beta$ $\gamma$cdots
$m_{\zeta/{\mathbb{F}}_{2}}(t)$ $t\kern-2.0pt+\kern-2.0pt1$ $t^{2}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{4}\kern-2.0pt+\kern-2.0ptt^{3}\kern-2.0pt+\kern-2.0pt1$ $t^{8}\kern-2.0pt+\kern-2.0ptt^{7}\kern-2.0pt+\kern-2.0ptt^{2}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{16}\kern-2.0pt+\kern-2.0ptt^{15}\kern-2.0pt+\kern-2.0ptt^{8}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$cdots
Proof.
Set $E={\mathbb{F}}_{2^{n}}$ and $F={\mathbb{F}}_{2}$. The polynomials in the last line of Table 4.8 are irreducible over $F$ (and hence can equal $m_{\zeta/F}(t)$), and $\zeta$ has order $12^{n_{2}}-$ 1. By Theorem 3.1, $x=R(y,z)$ is a root of $t^{2}+t+y$ where $z\in E$ is an element of degree $e=n_{2}$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Let $K$ be the subfield of $E$ with $|K:F|=n_{2}$ and $|E:K|=n_{2^{\prime}}$. We choose an element $\zeta\in K$ with $K={\mathbb{F}}_{2}(\zeta)$ and $\textup{Tr}^{K}_{F}(\zeta)=1$, and set $z=(n_{2^{\prime}})^{-1}\zeta$. If $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{j}}$, then $s_{0}=s_{2e}=0$ and $s_{i+2e}=s_{i}$, so it suffices to list $x_{i}=(n_{2^{\prime}})^{-1}s_{i}$ for $0\leqslant i<2e$. Indeed, since $s_{e}=1$ and $s_{i+e}=1+s_{i}$, it suffices to list $x_{i}$ for $0\leqslant i<e$.
(a) Suppose that $n_{2}=1$. Either set $\zeta=1$, $z=\frac{1}{n}$, or set $p=2$ in Corollary 4.2. Then $x^{2}+x+y=0$ where $x\kern-1.0pt=\kern-1.0pt\sum_{i=0}^{n-1}\frac{i}{n}y^{2^{i}}\kern-1.0pt=\kern-1.0pt\sum_{j=0}^{(n-1)/2}y^{2^{2j+1}}$ and $x_{i}=0,1$ for $i\equiv 0,1\pmod{2}$.
(b) Suppose that $n_{2}=2$ or $n\equiv 2\pmod{4}$. Set $\zeta=\omega$ or set $p=2$ into Corollary 4.6. Since $2\nmid(n/2)$ the value of $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r\omega\right)$ is $0,\omega,1,\omega^{2}$ when $i\equiv 0,1,2,3\pmod{4}$. The formulas for $x$ in cases (a) and (b) were given previously by Lahtonen [3].
(c) Suppose that $n_{2}=4$, i.e. $n\equiv 4\pmod{8}$. Then $\zeta=\alpha\in{\mathbb{F}}_{16}$ has order $12^{4}-$ 1 and satisfies $\textup{Tr}^{K}_{F}(\zeta)=1$. Hence $s_{i}=0,\alpha,\alpha^{13},\alpha^{6},1,\alpha^{12},\alpha^{7},\alpha^{8}$ for $i\equiv 0,1,\dots,7\pmod{8}$.
(d) Suppose that $n_{2}=8$, i.e. $n\equiv 8\pmod{16}$. Then $\zeta=\beta\in{\mathbb{F}}_{256}$ has order $12^{8}-$ 1 and satisfies $\textup{Tr}^{K}_{F}(\zeta)=1$. Hence $s_{i}=0,\beta,\beta^{100},\beta^{189},\beta^{29},\beta^{60},\beta^{154},\beta^{177}$ for $i\equiv 0,1,\dots,7\pmod{16}$. Also $s_{i+8}=1+s_{i}$ for $i\equiv 0,1,\dots,7\pmod{16}$.
(e) Suppose that $n_{2}=16$, i.e. $n\equiv 16\pmod{32}$. Then $\zeta=\gamma\in{\mathbb{F}}_{2^{16}}$ has order $12^{16}-$ 1 and $\textup{Tr}^{K}_{F}(\zeta)=1$. Hence $s_{i}=0,\gamma,\gamma^{64409},\gamma^{48754},\gamma^{27742},\gamma^{48469},\gamma^{1146},\gamma^{22404},\gamma^{64313},\gamma^{47682},$ $\gamma^{63219},\gamma^{45929},\gamma^{55680},\gamma^{46875},\gamma^{7495},\gamma^{32204}$ and $s_{i+16}=1+s_{i}$ for $i\equiv 0,1,\dots,15\pmod{32}$. ∎
Remark 4.9.
If $e=n_{p}$, then there are $p^{e-1}$ choices for an element $\zeta\in K={\mathbb{F}}_{p^{e}}$ with $\textup{Tr}^{K}_{F}(\zeta)=1$ where $F={\mathbb{F}}_{p}$. In Lemma 4.8 we chose $\zeta$ so that the minimal polynomial $m_{\zeta/F}(t)$ has many zero coefficients, and $\zeta$ has order $12^{e}-$ 1. However, there are many choices for $\zeta$ in general, and there appears to be no canonical choice.