### v2

# Hilbert’s Theorem 90, periodicity, and roots of Artin-Schreier polynomials

Abstract
Let $E/F$ be a cyclic field extension of degree $n$, and let $\sigma$ generate the group $\textup{Gal}(E/F)$. If $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}\sigma^{i}y=0$, then the additive form of Hilbert’s Theorem 90 asserts that $y=\sigma x-x$ for some $x\in E$. When $E$ has characteristic $p>0$ we prove that $x$ gives rise to a periodic sequence $x_{0},x_{1},\dots$ which has period $pn_{p}$, where $n_{p}$ is the largest $p$-power that divides $n$. We also show, if $y$ lies in the finite field ${\mathbb{F}}_{p^{n}}$, then the roots of a reducible Artin-Schreier polynomial $t^{p}-t-y$ have the form $x+u$ where $u\in{\mathbb{F}}_{p}$ and $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ for some $z\in{\mathbb{F}}_{p^{e}}$ with $e=n_{p}$. Furthermore, the sequence $\left(\sum_{j=0}^{i-1}z^{p^{j}}\right)_{i\geqslant 0}$ is periodic with period $pe$. Keywords: Hilbert’s Theorem 90, Artin-Schreier polynomials, periodic sequence 2020 Mathematics Subject Classification: 11T06, 12F10, 11T55
1 Introduction
Given a quadratic equation $ax^{2}+bx+c=0$ over a finite field ${\mathbb{F}}_{2^{n}}$ of order $2^{n}$, Chen [1] gives ‘formulas’ for the roots that involve $a,b,c$ and some element of trace 1. We give a unified treatment of these formulas which in prime characteristic $p$ depends on $n_{p}$, the largest $p$-power divisor of $n$. This is related the additive version of Hilbert’s Theorem 90 and to sequences of period $pn_{p}$. We say that a sequence $(x_{i})_{i\geqslant 0}$ has period $e$ if $x_{i}=x_{j}$ whenever $i\equiv j\pmod{e}$, and $e$ is minimal. In particular, we prove that a root of a reducible quadratic $x^{2}+x+y\in{\mathbb{F}}_{2^{n}}[x]$ has the form $x=\sum_{i=0}^{n-1}x_{i}y^{2^{i}}$, $x_{i}\in{\mathbb{F}}_{2^{n}}$, where the sequence $(x_{i})_{i\geqslant 0}$ has period $2n_{2}$. If $n_{2}=1$, we may take $x_{i}=0,1$ whenever $i\equiv 0,1\pmod{2}$, and if $n_{2}=2$, we may take $x_{i}=0,\omega,1,\omega^{2}$ whenever $i\equiv 0,1,2,3\pmod{4}$: here $\omega^{2}+\omega+1=0$ and $\omega$ is a primitive cube root of 1. The periodic sequence $(x_{i})_{i\geqslant 0}$ is described in detail, and for larger values of $n_{2}$, in Lemma 4.7.
This note proves the following two main results which we contextualize below. Given a finite Galois extension $E/F$ with group $\Gamma$, the trace map $E\to F$ is $\textup{Tr}^{E}_{F}(y)=\sum_{\gamma\in\Gamma}\gamma y$.
Theorem 1.1.
Let $E/F$ be a cyclic extension of degree $n$. Suppose $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $y,z\in E$ satisfy $\textup{Tr}^{E}_{F}(y)=0$, $\textup{Tr}^{E}_{F}(z)=1$, then $y=\sigma x-x$ where $x=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$.
The element $z$ in Theorem 1.1 satisfies additional properties in prime characteristic.
Theorem 1.2.
Let $E/F$ be a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\textup{char}(F)=p$. Let $n$ have $p$-part $n_{p}$, and $p^{\prime}$-part $n_{p^{\prime}}=n/n_{p}$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=0$, then
  1. (a)
    $F(z)$ contains the fixed field of $\sigma^{n_{p}}$, and $z$ may be chosen so that $|F(z):F|=n_{p}$;
  2. (b)
    the sequence $(\sum_{j=0}^{i-1}\sigma^{j}z)_{i\geqslant 0}$ is periodic with period $pe$ where $e=|F(z):F|$ and $n_{p}\mid e$.
Polynomials of the form $t^{p}-t-y$ where $y\in E$ were studied by Artin and Schreier in the context of generalizing Galois theory to characteristic $p$. Our theorems relate to roots of these Artin-Schreier polynomials. Suppose $\textup{char}(E)=p>0$ and the $p$th power map $\sigma\colon x\mapsto x^{p}$ of $E$ generates $\textup{Gal}(E/F)$, e.g. when $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$. If $\textup{Tr}^{E}_{F}(y)=0$ and $\textup{Tr}^{E}_{F}(z)=1$, then $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ is a root of the polynomial $t^{p}-t-y$. For example, if $n_{p}=1$, we may take $z=\frac{1}{n}$ and $x_{i}\coloneq\sum_{j=0}^{i-1}z^{p^{j}}=\frac{i}{n}$ defines a sequence of period $p$, and $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{p^{i}}$ is a root of $t^{p}-t-y$. Because different choices of $n_{p}$ involve different choices for $z$ with $\textup{Tr}^{E}_{F}(z)=1$, this suggests that there are infinitely many (polynomial) formulas for the roots of polynomials $t^{p}-t-y\in E[t]$. If $E={\mathbb{F}}_{p^{n}}$, $F={\mathbb{F}}_{p}$ and $e=n_{p}$, then finding $z$ in ${\mathbb{F}}_{p^{e}}$ with $\textup{Tr}^{E}_{F}(z)=1$ is easy in practise: with probability $1-\frac{1}{p}$ a random element $\zeta\in{\mathbb{F}}_{p^{e}}$ has $\tau\coloneq\textup{Tr}^{E}_{F}(\zeta)\neq 0$, so $z=\tau^{-1}\zeta$ has $\textup{Tr}^{E}_{F}(z)=1$.
There are two versions of Hilbert’s Theorem 90: the additive form and the multiplicative form: we abbreviate these AH90 and MH90, respectively. Theorem 1.1 is different from usual version of AH90 in that the formula for $x$ is very slightly different. This difference leads to Theorem 1.2. We remark first that MH90 is so numbered because it appears as Satz 90 in [2] on p. 272, and Hilbert never actually proved AH90. The standard proof of AH90 is given in Lang [5, p. 215] and it follows Emil Artin’s ideas. In Theorem 1.1, $x$ equals $R(y,z)\coloneq\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\sigma^{j}z\sigma^{i}y$ where $z\in E$ satisfies $\textup{Tr}^{E}_{F}(z)=1$ and $y=\sigma x-x$ holds. By contrast, the standard proof of AH90 uses a different element, namely $x^{\prime}=R(z,y)$ and $y=x^{\prime}-\sigma x^{\prime}$ holds. The proofs of AH90 I have found, assume $\textup{Tr}^{E}_{F}(z)\neq 0$ (but renormalizing to get $\textup{Tr}^{E}_{F}(z)=1$ is easy) and show that
$x^{\prime}$ $=\frac{1}{\textup{Tr}^{E}_{F}(z)}\left(y\sigma(z)+(y+\sigma(y))\sigma^{2}(z)+\cdots+(y+\sigma(y)+\cdots+\sigma^{n-2}(y))\sigma^{n-1}(z)\right)$
$=\frac{1}{\textup{Tr}^{E}_{F}(z)}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\sigma^{j}y\sigma^{i}z=\frac{1}{\textup{Tr}^{E}_{F}(z)}R(z,y)$
satisfies $y=x^{\prime}-\sigma x^{\prime}$. The point here is that writing $x$ in the form $\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ for some $x_{i}\in E$, suggests studying the sequence $(x_{i})_{i\geqslant 0}$, whereas the form of $x^{\prime}$ does not.
We remark that the expression $x^{\prime}=R(z,y)$ is very natural in the context of Galois cohomology. Indeed, this form suggests itself via Emmy Noether’s equations, and these equations lead to analogous proofs of AH90 and MH90 as we explain. Emmy Noether generalized the method used to prove AH90 and MH90 by introducing 1-cocyles, to handle a noncyclic (finite) Galois group $\Gamma\coloneq\textup{Gal}(E/F)$, see [3, Theorems 4.29 and 4.30]. Her result [3, Theorem 4.29] says: given a function $y\colon\Gamma\to E\setminus\{0\}$ satisfying $y_{\beta\gamma}=y_{\beta}\beta(y_{\gamma})$ for $\beta,\gamma\in\Gamma$, then for some $z\in E$, $x=\sum_{\beta\in\Gamma}y_{\beta}\beta(z)$ satisfies $y_{\gamma}=x\gamma(x)^{-1}$ for all $\gamma\in\Gamma$. Her result [3, Theorem 4.30] says: given a function $y\colon\Gamma\to E$ satisfying $y_{\beta\gamma}=y_{\beta}+\beta(y_{\gamma})$ for $\beta,\gamma\in\Gamma$, then for some $z\in E$, $x=\textup{Tr}^{E}_{F}(z)^{-1}\sum_{\beta\in\Gamma}y_{\beta}\beta(z)$ satisfies $y_{\gamma}=x-\gamma(x)$ for all $\gamma\in\Gamma$. When $\Gamma=\langle\sigma\rangle$ is cyclic of order $n$, the 1-cocycle conditions imply that $y_{\sigma^{i}}=y(\sigma y)\cdots(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$, and similarly in the additive case $y_{\sigma^{i}}=y+(\sigma y)+\cdots+(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$. This gives rise to the multiplicative formula $x=\sum_{i=0}^{n-1}\left(\prod_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$ and when $\textup{Tr}^{E}_{F}(z)=1$, the additive formula $x=R(z,y)\coloneq\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$, respectively. Theorem 1.1 uses the related expression $R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$. Lemma 3.1 in some sense measures the asymmetry between $R(y,z)$ and $R(z,y)$ via the equation $R(y,z)+R(z,y)=-\sum_{i=0}^{n-1}\sigma^{i}(yz)$.
2 Background
This work fits into a very classical setting, namely generalizing Galois theory to characteristic $p$. The following important result follows from the AH90, see [3, Theorem 4.33] and [5, VIII, Theorem 11] for proofs of Theorem 2.1.
Theorem 2.1 (Artin-Schreier).
Let $E$ be field of prime characteristic $p$.
  1. (a)
    Let $E/F$ be a cyclic extension of degree $p$. Then $E=F(x)$ where $x^{p}-x\in F\setminus\{0\}$.
  2. (b)
    Given $y\in E$, the polynomial $t^{p}-t-y$ is either irreducible over $E$, or it has $p$ distinct roots $x+u$ where $x\in E$ and $u\in{\mathbb{F}}_{p}$.
The trace map $\textup{Tr}_{F}^{E}(x)=\sum_{\gamma\in\textup{Gal}(E/F)}\gamma(x)$ of a finite Galois extension $E/F$, is an $F$-linear map $E\to F$ which is nonzero by Artin’s Lemma [5, VIII, Theorem 7]. In certain cases, the trace map can be used to distinguish whether or not $t^{p}-t+y$ is irreducible.
Lemma 2.2.
Let $E/F$ be a cyclic Galois extension where $|E:F|=\textup{char}(F)=p$. Suppose that $\textup{Gal}(E/F)=\langle\sigma\rangle$ where $\sigma\colon E\to E$ is the $p$th power map $\sigma(x)=x^{p}$. For $y\in E$, the polynomial $t^{p}-t+y$ is reducible over $E$ if and only if $\textup{Tr}_{F}^{E}(y)=\sum_{i=0}^{p-1}y^{p^{i}}=0$. Furthermore, $E=F(z)$ for some $z\in E$ where $z^{p}-z^{p-1}+1=0$ and $\textup{Tr}^{E}_{F}(z)=1$.
Proof.
If $\textup{Tr}_{F}^{E}(y)=0$, then $y=z-\sigma(z)=z-z^{p}$ for some $z\in E$ by AH90. Hence $t^{p}-t+y$ is reducible over $E$. Conversely, suppose if $t^{p}-t+y$ is reducible over $E$, then $t^{p}-t+y$ has a root $z\in E$, by Theorem 2.1(b). Using $\sigma z=z^{p}$ shows
$0=\textup{Tr}^{E}_{F}(0)=\textup{Tr}^{E}_{F}(z^{p}-z+y)=\textup{Tr}^{E}_{F}(\sigma z)-\textup{Tr}^{E}_{F}(z)+\textup{Tr}^{E}_{F}(y)=\textup{Tr}^{E}_{F}(y).$
Finally, set $y=1$. Then $\textup{Tr}_{F}^{E}(1)=p\cdot 1=0$ and some $\zeta\in E$ satisfies $t^{p}-t+1=0$. However, $\sigma\zeta=\zeta^{p}=\zeta-1$ implies $\sigma^{i}\zeta=\zeta-i$ so $\zeta$ has $p$ distinct conjugates in $E$ and $E=F(\zeta)$. Therefore $E=F(z)$ where $z=\zeta^{-1}$ is a root of the (reverse) polynomial $t^{p}(t^{-p}-t^{-1}+1)=t^{p}-t^{p-1}+1$. Since the conjugates $z,\sigma z,\dots,\sigma^{p-1}z$ are pairwise distinct, $t^{p}-t^{p-1}+1$ is also irreducible over $F$. The coefficient of $t^{p-1}$ shows that $\textup{Tr}^{E}_{F}(z)=1$. ∎
We note that only part of Theorem 2.1(b) generalizes. If ${\mathbb{F}}_{q}\leqslant F\leqslant E$, and $t^{q}-t-y$ has a root $x\in E$, then $t^{q}-t-y$ factors as $\prod_{a\in{\mathbb{F}}_{q}}(t-(x+a))$. However, if $t^{q}-t-y$ has no root in $E$, we can not conclude that $t^{q}-t-y$ is irreducible over $E$. For example, if $q=4$ then $t^{4}+t+1$ factors as $(t^{2}+t+\omega)(t^{2}+t+\omega^{2})$ over ${\mathbb{F}}_{4}={\mathbb{F}}_{2}(\omega)$.
[…]
3 Periodicity and Hilbert’s additive Theorem 90
Let $E/F$ be a Galois extension with group $\Gamma=\textup{Gal}(E/F)$. Thus $\textup{Tr}^{E}_{F}(\sigma x)=\textup{Tr}^{E}_{F}(x)$ for $\sigma\in\Gamma$. Hence if $y\in E$ has the form $\sigma x-x$ for some $x\in E$ and $\sigma\in\Gamma$, then $\textup{Tr}^{E}_{F}(y)=0$. If $\Gamma$ is finite and cyclic, then the converse holds by AH90. In this section $E/F$ is a cyclic extension of degree $n$, and $\textup{Gal}(E/F)=\langle\sigma\rangle$. Thus $\sigma^{n}=1$ and $\textup{Tr}_{F}^{E}(y)=\sum_{i=0}^{n-1}\sigma^{i}y$.
Proof of Theorem 1.1.
Suppose that $x$ has the form $\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ where $x_{0},\dots,x_{n-1}\in E$. We aim to solve for the coefficients $x_{i}$. Observe that
$\sigma x-x=-x_{0}y+\sum_{i=1}^{n-1}(\sigma x_{i-1}-x_{i})\sigma^{i}y+\sigma x_{n-1}\sigma^{n}y.$
Assume for the moment that there exists an element $z\in E$ satisfying
(1) $\sigma x_{i-1}-x_{i}=-z\quad\textup{for $1\leqslant i<n$, and}\quad\sigma x_{n-1}-x_{0}=1-z.$
Substituting these equations and using $\sigma^{n}y=y$ gives
$\sigma x-x=-x_{0}y-z\sum_{i=1}^{n-1}\sigma^{i}y+(1-z+x_{0})y=y-z\sum_{i=0}^{n-1}\sigma^{i}y=y$
as desired. Next, assume that $z$ and $x_{0}$ are given. The equations $\sigma x_{i-1}-x_{i}=-z$ show that $x_{i}=\sigma^{i}x_{0}+\sigma^{i-1}z+\cdots+\sigma z+z$ holds for $1\leqslant i<n$. Rewrite the $n$ equations (1) as $\sigma^{n-i+1}x_{i-1}-\sigma^{n-i}x_{i}=-\sigma^{n-i}z$ for $1\leqslant i<n$ and $\sigma x_{n-1}-x_{0}=1-z$. Adding these equations gives $0=\sigma^{n}x_{0}-x_{0}=1-\textup{Tr}_{F}^{E}(z)$. Hence $\textup{Tr}_{F}^{E}(z)=1$ must hold.
To see that these necessary equations are sufficient, choose $z\in E$ so that $\textup{Tr}_{F}^{E}(z)=1$ which is possible as the trace map $\textup{Tr}_{F}^{E}\colon E\to F$ is surjective. Next, define $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where the empty sum $x_{0}=\sum_{j=1}^{0}\sigma^{j}z$ is 00 by convention. Then the $n$ equations in (1) hold, and $\sigma x-x$ equals $y$, as claimed. ∎
Lemma 3.1.
Let $E/F$ be a cyclic Galois extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$. If $y,z\in E$ have traces $\textup{Tr}^{E}_{F}(y)=0$ and $\textup{Tr}^{E}_{F}(z)=1$, then $R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$ satisfies $R(y,z)=-\textup{Tr}^{E}_{F}(yz)-R(z,y)$, $\sigma R(y,z)-R(y,z)=y$ and $R(z,y)-\sigma R(z,y)=y$.
Proof.
The equations $\sum_{i=0}^{n-1}\sigma^{i}y=0$ and $\sum_{j=0}^{n-1}\sigma^{j}z=1$ imply
$R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y=\sum_{i=0}^{n-1}\left(1-\sum_{j=i}^{n-1}\sigma^{j}z\right)\sigma^{i}y=-\sum_{i=0}^{n-1}\left(\sum_{j=i}^{n-1}\sigma^{j}z\right)\sigma^{i}y.$
Swapping the order of summation, swapping $i\leftrightarrow j$, and noting $\textup{Tr}^{E}_{F}(zy)=\sum_{i=0}^{n-1}\sigma^{i}(zy)$ gives
$R(y,z)=-\sum_{j=0}^{n-1}\sum_{i=0}^{j}\sigma^{j}z\sigma^{i}y=-\sum_{i=0}^{n-1}\sum_{j=0}^{i}\sigma^{i}z\sigma^{j}y=-\textup{Tr}^{E}_{F}(zy)-\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z.$
Therefore $R(y,z)=-\textup{Tr}^{E}_{F}(yz)-R(z,y)$. Theorem 1.2 proves that $\sigma R(y,z)-R(y,z)=y$ holds. Since $\sigma\textup{Tr}^{E}_{F}(yz)=\textup{Tr}^{E}_{F}(yz)$ it follows that $R(z,y)-\sigma R(z,y)=y$, as claimed. ∎
We prove in Theorem 1.2 that when $\textup{char}(F)=p>0$, the element $x$ in Theorem 1.1 has an intrinsic periodicity. This observation is used in Corollary 4.1 to give a polynomial formula for the roots of (reducible) Artin-Schreier polynomials.
Proof of Theorem 1.2.
(a) First, we show that the fixed subfield $K$ of $\sigma^{n_{p}}$ contains an element $z$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Since $n_{p^{\prime}}=n/n_{p}\neq 0$ in $F$ and the map $\textup{Tr}^{K}_{F}\colon K\to F$ is surjective, there exists an element $z\in K$ such that $\textup{Tr}^{K}_{F}(z)=(n_{p^{\prime}})^{-1}$. Therefore $\textup{Tr}^{E}_{K}(z)=|E:K|z=n_{p^{\prime}}z$ and hence $\textup{Tr}^{E}_{F}(z)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(z))=n_{p^{\prime}}\textup{Tr}^{K}_{F}(z)=1$. Next, suppose that $L=F(z)$ and $\textup{Tr}^{E}_{F}(z)=1$. Then $1=\textup{Tr}^{L}_{F}(\textup{Tr}^{E}_{L}(z))=|E:L|\textup{Tr}^{L}_{F}(z)$. Therefore $p\nmid|E:L|$, so $n_{p}=|K:F|$ divides $e=|L:F|$ as claimed. In summary, the minimal polynomial over $F$ of an element $z$ with $\textup{Tr}^{E}_{F}(z)=1$ has degree a multiple of $n_{p}$, and degree precisely $n_{p}$ if $z$ lies in $K\setminus K_{0}$ where $K_{0}$ is the unique (maximal) subfield of $K$ with $F\leqslant K_{0}<K$ and $|K:K_{0}|=p$.
(b) Let $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where $\textup{Tr}^{E}_{F}(z)=1$. Let $L=F(z)$ and $e=|L:F|$. Hence $x_{e}=\textup{Tr}^{L}_{F}(z)\neq 0$ by part (b) and $\sigma^{j}z=z$ implies $e\mid j$. Suppose that $x_{k}=0$ where $0\leqslant k<pe$ and write $k=ie+j$ where $0\leqslant i<p$ and $0\leqslant j<e$. The equation $x_{ie+j}=ix_{e}+x_{j}$ shows that $x_{0}=x_{pe}=0$. Thus $x_{k}=ix_{e}+x_{j}=0$ and $x_{e}\in F$ gives
$x_{j+1}-z=\sigma x_{j}=\sigma(-ix_{e})=-ix_{e}=x_{j}.$
Whence $\sigma^{j}z=x_{j+1}-x_{j}=z$. This implies that $e\mid j$ and hence $j=0$. However, $0=x_{k}=x_{ie}=ix_{e}$ shows that $0=i=j=k$. Therefore $x_{\ell}\neq 0$ holds for $0<\ell<pe$, and the sequence $x_{0},x_{1},\dots$ is periodic, with period precisely $pe$. ∎
4 Formulas for the roots of Artin-Schreier polynomials
In this section we assume that $F$ has prime characteristic $p$. Set $q=p^{f}$. Two examples with $\textup{Gal}(E/F)=\langle\sigma\rangle$ of order $n$ and $\sigma(x)=x^{q}$ for $x\in E$ are the finite fields $E={\mathbb{F}}_{q^{n}}$, $F={\mathbb{F}}_{q}$, and the function fields $E={\mathbb{F}}_{q^{n}}(u)$, $F={\mathbb{F}}_{q}(u)$. We shall give a number of different applications of Theorem 1.2 depending on the value of $n_{p}$. Observe first that Theorem 1.1 gives a polynomial formula for $x$ when $\sigma(x)=x^{q}$ as noted below.
Corollary 4.1.
Suppose that $\textup{char}(F)=p$, $q=p^{f}$, and $E/F$ is a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=0$, then $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{q^{j}}y^{q^{i}}$ satisfies $x^{q}-x-y=0$.
When $n$ is coprime to $p$, i.e. $n_{p}=1$, the formula for $x$ in Corollary 4.1 simplifies.
Corollary 4.2.
Suppose that ${\mathbb{F}}_{q}\leqslant F$, $q=p^{f}$, and $E/F$ is a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$. If $y\in E$, $\textup{Tr}^{E}_{F}(y)=0$ and $n_{p}=1$, then $t^{q}-t-y=\prod_{u\in{\mathbb{F}}_{q}}(t-x-u)$ where $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$.
Proof.
By Theorem 1.2(b), we may choose $z$ to lie in $F$ and so $1=\textup{Tr}^{E}_{F}(z)=nz$. Hence $z=\frac{1}{n}$ and $x=R(y,z)=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$ satisfies $x^{q}-x-y=0$ by Corollary 4.1. Therefore $x+u$ is a root of $t^{q}-t-y$ for each $u\in{\mathbb{F}}_{q}\leqslant F$, and the claimed factorization follows.∎
Remark 4.3.
When $\textup{char}(F)=p\nmid n$, Corollary 4.2 implies that all the roots of $t^{q}-t-y$ lie in the ${\mathbb{F}}_{p}$-subspace $\langle y,y^{q},\dots,y^{q^{n-1}}\rangle$ of ${\mathbb{F}}_{q^{n}}$. Since $\sum_{i=0}^{n-1}y^{q^{i}}=0$, the roots of $t^{q}-t-y$ lie in the (potentially larger) ${\mathbb{F}}_{q}$-subspace $\{w\in E\mid\textup{Tr}^{E}_{F}(w)=0\}$ of $E$ of codimension 1.
When $n_{p}=p$, i.e. $p\mid n$ and $p^{2}\nmid n$, the element $z$ in Corollary 4.1 can be chosen to lie in the fixed field of $\sigma^{p}$, and to have a specific minimal polynomial.
Corollary 4.4.
Let $E/F$ be a cyclic field extension of degree $n$ where $n_{p}=\textup{char}(F)=p$. Suppose that $\textup{Gal}(E/F)=\langle\sigma\rangle$ where $\sigma\colon E\to E$ is the $p$th power map $\sigma(x)=x^{p}$. If $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}y^{p^{i}}=0$, and $\zeta$ satisfies $\zeta^{p}-\zeta^{p-1}+1=0$. Then $x^{p}-x-y=0$ where $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$ and $\zeta^{p^{p}-1}=1$.
Proof.
It follows from $n_{p}=p$, that $n_{p^{\prime}}=n/p$ is coprime to $p$. Let $K$ be the fixed field of $\sigma^{p}$. Then $K$ contains an element $\zeta$ satisfying $\zeta^{p}-\zeta^{p-1}+1=0$ and $\textup{Tr}^{L}_{F}(\zeta)=1$ by Lemma 2.2. Therefore $\textup{Tr}^{E}_{F}(\zeta)=n/p$ and $\textup{Tr}^{E}_{F}(z)=1$ where $z=(n/p)^{-1}\zeta$. By Theorem 1.1, $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$ satisfies $x^{p}-x-y=0$. Finally, $\sigma(\zeta)=\zeta^{p}$ implies that $\zeta=\sigma^{p}(\zeta)=\zeta^{p^{p}}$, so that $\zeta^{p^{p}-1}=1$ as $\zeta\neq 1$. ∎
We shall now apply Corollary 4.1 in the case that $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$, so that $E/F$ is cyclic and $\textup{Gal}(E/F)=\langle\sigma\rangle$ has order $n$ where $\sigma(x)=x^{p}$ is the $p$th power map.
Let $p,r$ be distinct primes. Denote by $\textup{ord}_{r}(p)$ the smallest positive integer $e$ such that $p^{e}\equiv 1\pmod{r}$ and note that $e\mid(r-1)$. The cyclotomic polynomial $\Phi_{r}(t)=\sum_{i=0}^{r-1}t^{i}$ of degree $r-1$ factors over ${\mathbb{F}}_{p}$ as a product of $(r-1)/e$ irreducible polynomials each of degree $e=\textup{ord}_{r}(p)$ by [6, Theorem 2.47].
Theorem 1.2 shows that $z$ may be chosen to lie in a subfield of order $p^{f}$ where $f={n_{p}}$. Hence $z^{p^{f}-1}=1$. Lemma 4.5 proves that, in many cases, $z$ may be chosen to have prime order $r\neq p$. This choice of $z$, however, may increase the period of the sequence $(x_{i})_{i\geqslant 0}$.
Lemma 4.5.
Let $p,r$ be unequal primes. Let $e\coloneq\textup{ord}_{r}(p)$ divide $n$. Set $E={\mathbb{F}}_{p^{n}}$, $K={\mathbb{F}}_{p^{e}}$ and $F={\mathbb{F}}_{p}$. There exists a primitive $r$th root of unity $\zeta\in K$ for which $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. If $p\nmid\frac{n}{e}$ and $y\in E$ has $\textup{Tr}^{E}_{F}(y)=0$, then $x=(\frac{n\tau}{e})^{-1}\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\zeta^{p^{j}}\right)y^{p^{i}}$ satisfies $x^{p}-x-y=0$. Finally, the sequence $(\sum_{j=0}^{i-1}\zeta^{p^{j}})_{i\geqslant 0}$ is periodic with period $ep$ where $n_{p}\mid e$.
Proof.
First, $r$ divides $p^{d}-1$ if and only if $e\mid d$. As $e\mid n$, all primitive $r$th roots of unity lie in $K$, and $F\leqslant K\leqslant E$ holds. The cyclotomic polynomial $\Phi_{r}(t)$ factors over $F$ as a product of $(r-1)/e$ irreducible polynomials each of degree $e$, so the coefficient of $t^{e-1}$ must be nonzero in at least one irreducible. Hence some $r$th root of unity $\zeta\in K$ has $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. Thus $\textup{Tr}^{E}_{K}(\zeta)=\frac{n}{e}\zeta$ and $\textup{Tr}^{E}_{F}(\zeta)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(\zeta))=\frac{n}{e}\textup{Tr}^{K}_{F}(\zeta)=\frac{n\tau}{e}\neq 0$. Therefore $z=(\frac{n\tau}{e})^{-1}\zeta\in K^{\times}$ satisfies $\textup{Tr}^{E}_{F}(z)=1$. The formula for $x=R(y,z)$ follows from Theorem 1.1 and the periodicity $ep$ follows from Theorem 1.2(b). ∎
Corollary 4.6.
Let $p$ be a prime where $p\equiv 2\pmod{3}$. Let $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$ where $n$ is even and let $\omega\in E$ satisfy $\omega^{2}+\omega+1=0$. If $p\nmid(n/2)$, and $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=0$, then the sum $x=(\frac{n}{2})^{-1}\sum_{i=0}^{n-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)y^{p^{i}}$ satisfies $x^{p}-x-y=0$ where $r_{i}\in\{0,1\}$ and $i\equiv r_{i}\pmod{2}$. Finally, the sequence $x_{0},x_{1},\dots$ is periodic of period $2p$ where $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)$ and $r_{i}=i\mod 2\in\{0,1\}$.
Proof.
Since $n$ is even and $p^{n}\equiv 1\pmod{3}$, $E$ contains a primitive cube root $\omega$ of 1. The condition $p\equiv 2\pmod{3}$ is equivalent to $\textup{ord}_{3}(p)=2$. Thus we may use Lemma 4.5 with $r=3$ and $\zeta=\omega$. However, $\omega^{p^{k}}$ is $\omega$ if $k$ is even, and $\omega^{2}$ if $k$ is odd. If $i=0,1,2,3,4,5,6$, then $s_{i}=\sum_{k=0}^{i-1}\omega^{p^{k}}$ equals 00, $\omega$, $\omega+\omega^{2}=-1$, $-1+\omega$, $-2$, $-2+\omega$, $-3$, etc. Therefore $s_{i}=-(i-r_{i})/2+r_{i}\omega$ where $r_{i}\equiv i\pmod{2}\in\{0,1\}$ and $r_{i}\in\{0,1\}$. The values of $x_{i}=(\frac{n}{2})^{-1}\left(\frac{i-r_{i}}{2}-r_{i}\omega\right)$ are periodic of period $2p$ by Theorem 1.2(b). Finally, $\frac{i-r_{i}}{2}=\lfloor\frac{i}{2}\rfloor$. ∎
The following result gives different quadratic ‘formulas’ over the finite field ${\mathbb{F}}_{2^{n}}$ that depend on the 2-part $n_{2}$ of $n$ and the choice of an element $z\in{\mathbb{F}}_{2^{n}}$ with absolute trace 1. The condition $n_{2}=2^{k}$ is equivalent to $n\equiv 2^{k}\pmod{2^{k+1}}$.
Lemma 4.7.
Let $y\in{\mathbb{F}}_{2^{n}}$ where $\sum_{i=0}^{n-1}y^{2^{i}}=0$. Then $x=\sum_{i=0}^{n-1}x_{i}y^{2^{i}}$ is a root of $t^{2}+t+y$ where $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ and $z\in{\mathbb{F}}_{2^{n}}$ has absolute trace $\sum_{i=0}^{n-1}z^{2^{i}}=1$ and has degree $n_{2}$. Write $z=(n/n_{2})^{-1}\zeta$. Table 4.7 lists choices for $\zeta$ and the minimal polynomial $m_{\zeta/{\mathbb{F}}_{2}}(t)$ of $\zeta$, as a function of the value of $n_{2}$.
$n_{2}$ $1$ $2$ $4$ $8$ $16$cdots
$\zeta$ $1$ $\omega$ $\alpha$ $\beta$ $\gamma$cdots
$m_{\zeta/{\mathbb{F}}_{2}}(t)$ $t\kern-2.0pt+\kern-2.0pt1$ $t^{2}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{4}\kern-2.0pt+\kern-2.0ptt^{3}\kern-2.0pt+\kern-2.0pt1$ $t^{8}\kern-2.0pt+\kern-2.0ptt^{7}\kern-2.0pt+\kern-2.0ptt^{2}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{16}\kern-2.0pt+\kern-2.0ptt^{15}\kern-2.0pt+\kern-2.0ptt^{8}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$cdots
Proof.
Set $E={\mathbb{F}}_{2^{n}}$ and $F={\mathbb{F}}_{2}$. The polynomials in the last line of Table 4.7 are irreducible over $F$ (and hence can equal $m_{\zeta/F}(t)$), and $\zeta$ has order $12^{n_{2}}-$ 1. By Theorem 1.1, $x=R(y,z)$ is a root of $t^{2}+t+y$ where $z\in E$ is an element of degree $e=n_{2}$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Let $K$ be the subfield of $E$ with $|K:F|=n_{2}$ and $|E:K|=n_{2^{\prime}}$. We choose an element $\zeta\in K$ with $K={\mathbb{F}}_{2}(\zeta)$ and $\textup{Tr}^{K}_{F}(\zeta)=1$, and set $z=(n_{2^{\prime}})^{-1}\zeta$. Let $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{j}}$. Then $s_{0}=s_{2e}=0$ and $s_{i+2e}=s_{i}$, so it suffices to list $x_{i}=(n_{2^{\prime}})^{-1}s_{i}$ for $0\leqslant i<2e$. Indeed, since $s_{e}=1$ and $s_{i+e}=1+s_{i}$, it suffices to list $x_{i}$ for $0\leqslant i<e$.
(a) Suppose that $n_{2}=1$. Either set $\zeta=1$, $z=\frac{1}{n}$, or set $p=2$ in Corollary 4.2. Then $x^{2}+x+y=0$ where $x\kern-1.0pt=\kern-1.0pt\sum_{i=0}^{n-1}\frac{i}{n}y^{2^{i}}\kern-1.0pt=\kern-1.0pt\sum_{j=0}^{(n-1)/2}y^{2^{2j+1}}$ and $x_{i}=0,1$ for $i\equiv 0,1\pmod{2}$.
(b) Suppose that $n_{2}=2$. Set $\zeta=\omega$ or set $p=2$ into Corollary 4.6. Since $2\nmid(n/2)$ the value of $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)$ is $0,\omega,1,\omega^{2}$ when $i\equiv 0,1,2,3\pmod{4}$. (Recall that $r_{i}=i-2\lfloor\frac{i}{2}\rfloor\in\{0,1\}$.) Equivalent formulas for $x$ in cases (a) and (b) were given previously by Chen [1, Theorems 1, 2] and by Lahtonen [4].
(c) Suppose that $n_{2}=4$. Then $\zeta=\alpha\in{\mathbb{F}}_{16}$ has order $12^{4}-$ 1 and $\textup{Tr}^{K}_{F}(\zeta)=1$. Further, the values of $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{i}}$ are $0,\alpha,\alpha^{13},\alpha^{6},1,\alpha^{12},\alpha^{7},\alpha^{8}$ for $i\equiv 0,1,\dots,7\pmod{8}$.
(d) Suppose that $n_{2}=8$. Then $\zeta=\beta\in{\mathbb{F}}_{256}$ has order $12^{8}-$ 1 and satisfies $\textup{Tr}^{K}_{F}(\zeta)=1$. Furthermore, the values of $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{i}}$ are $0,\beta,\beta^{100},\beta^{189},\beta^{29},\beta^{60},\beta^{154},\beta^{177}$ for $i\equiv 0,1,\dots,7\pmod{16}$. Also $s_{i+8}=1+s_{i}$ for $i\equiv 0,1,\dots,7\pmod{16}$.
(e) Suppose that $n_{2}=16$. Then $\zeta=\gamma\in{\mathbb{F}}_{2^{16}}$ has order $12^{16}-$ 1 and $\textup{Tr}^{K}_{F}(\zeta)=1$. In addition, $s_{i}=\sum_{j=0}^{i-1}\zeta^{2^{j}}$ equals $0,\gamma,\gamma^{64409},\gamma^{48754},\gamma^{27742},\gamma^{48469},\gamma^{1146},\gamma^{22404},\gamma^{64313},\gamma^{47682},$ $\gamma^{63219},\gamma^{45929},\gamma^{55680},\gamma^{46875},\gamma^{7495},\gamma^{32204}$ and $s_{i+16}=1+s_{i}$ for $i\equiv 0,1,\dots,15\pmod{32}$. ∎
Remark 4.8.
If $e=n_{p}$ and $F={\mathbb{F}}_{p}$, then there are $p^{e-1}$ choices for an element $\zeta\in K={\mathbb{F}}_{p^{e}}$ with $\textup{Tr}^{K}_{F}(\zeta)=1$. In Lemma 4.7 we chose $\zeta$ so that (a) $\sum_{i=0}^{n_{2}-1}\zeta^{2^{i}}=1$, (b) the minimal polynomial $m_{\zeta/F}(t)$ has many zero coefficients, and (c) $\zeta$ has multiplicative order $12^{e}-$ 1. There are many choices for $\zeta$ with these three properties, and there appears to be no canonical choice in general.