### v3

# Hilbert’s Theorem 90, periodicity, and roots of Artin-Schreier polynomials

Abstract
Let $E/F$ be a cyclic field extension of degree $n$, and let $\sigma$ generate the group $\textup{Gal}(E/F)$. If $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}\sigma^{i}y=0$, then the additive form of Hilbert’s Theorem 90 asserts that $y=\sigma x-x$ for some $x\in E$. When $E$ has characteristic $p>0$ we prove that $x$ gives rise to a periodic sequence $x_{0},x_{1},\dots$ which has period $pn_{p}$, where $n_{p}$ is the largest $p$-power that divides $n$. We also show, if $y$ lies in the finite field ${\mathbb{F}}_{p^{n}}$, then the roots of a reducible Artin-Schreier polynomial $t^{p}-t-y$ have the form $x+u$ where $u\in{\mathbb{F}}_{p}$ and $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ for some $z\in{\mathbb{F}}_{p^{e}}$ with $e=n_{p}$. Furthermore, the sequence $\left(\sum_{j=0}^{i-1}z^{p^{j}}\right)_{i\geqslant 0}$ is periodic with period $pe$.
Keywords: Hilbert’s Theorem 90, Artin-Schreier polynomials, periodic sequences
2020 Mathematics Subject Classification: 11T06, 12F10, 11T55
1 Introduction
Galois was initially concerned to determine when the roots of polynomial $f(x)$ over a field $F$ of characteristic zero could be expressed as a formula involving the symbols $+,-,\times,/$ and $n$th roots for $n\geqslant 2$. If $F$ has prime characteristic, then the formulas were adapted to included roots of Artin-Schreier polynomials. This note explores different formulas for the roots, replacing roots of Artin-Schreier polynomials with polynomial expressions.
Given a quadratic equation $ax^{2}+bx+c=0$ over a finite field ${\mathbb{F}}_{2^{n}}$ of order $2^{n}$, Chen [1] gives ‘formulas’ for the roots that involve $a,b,c$ and some element of trace 1. We give a unified treatment of these formulas, which in prime characteristic $p$ depends on $n_{p}$, the largest $p$-power divisor of $n$. This is related to the additive version of Hilbert’s Theorem 90 and to sequences of period $pn_{p}$. We say that a sequence $(x_{i})_{i\geqslant 0}$ has period $e$ if $x_{i}=x_{j}$ whenever $i\equiv j\pmod{e}$, and $e$ is minimal. We prove that a root of a reducible polynomial $x^{p}-x-y\in{\mathbb{F}}_{p^{n}}[x]$ has the form $x=\sum_{i=0}^{n-1}x_{i}y^{p^{i}}$, $x_{i}\in{\mathbb{F}}_{p^{n}}$, where the sequence $(x_{i})_{i\geqslant 0}$ has period $pn_{p}$. In the case that $p=2$ and $n$ is odd, we may take $x_{i}=i\in\{0,1\}$, and if $p=2$ and $n$ is twice odd, we may take $x_{i}=0,\omega,1,\omega^{2}$ whenever $i\equiv 0,1,2,3\pmod{4}$: here $\omega$ is a primitive cube root of 1. The periodic sequence $(x_{i})_{i\geqslant 0}$ is described in detail, for larger values of the 2-part $n_{2}$ of $n$, in Lemma 4.9.
This note proves the following two main results which we contextualize below. Given a finite Galois extension $E/F$ with group $\Gamma$, the trace map $E\to F$ is $\textup{Tr}^{E}_{F}(y)=\sum_{\gamma\in\Gamma}\gamma y$.
Theorem 1.1.
Let $E/F$ be a cyclic extension of degree $n$. Suppose $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $y,z\in E$ satisfy $\textup{Tr}^{E}_{F}(y)=0$, $\textup{Tr}^{E}_{F}(z)=1$, then $y=\sigma x-x$ where $x=\sum_{i=0}^{n-1}\kern-1.1pt\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\kern-1.1pt\sigma^{i}y$.
The element $z$ in Theorem 1.1 satisfies additional properties in prime characteristic.
Theorem 1.2.
Let $E/F$ be a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\textup{char}(F)=p$. Let $n$ have $p$-part $n_{p}$, so $n_{p}\mid n$ and $pn_{p}\nmid n$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=0$, then
  1. (a)
    $F(z)$ contains the fixed field of $\sigma^{n_{p}}$, and $z$ may be chosen so that $|F(z):F|=n_{p}$;
  2. (b)
    the sequence $(\sum_{j=0}^{i-1}\sigma^{j}z)_{i\geqslant 0}$ is periodic with period $pe$ where $e=|F(z):F|$ and $n_{p}\mid e$.
Polynomials of the form $t^{p}-t-y$ where $y\in E$ were studied by Artin and Schreier in the context of generalizing Galois theory to characteristic $p$: they are analogues of $p$th roots in characteristic 0. Our theorems relate to roots of these Artin-Schreier polynomials. Suppose $\textup{char}(E)=p>0$ and the $p$th power map $\sigma\colon x\mapsto x^{p}$ of $E$ generates $\textup{Gal}(E/F)$, e.g. when $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$. If $\textup{Tr}^{E}_{F}(y)=0$ and $\textup{Tr}^{E}_{F}(z)=1$, then $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ is a root of the polynomial $t^{p}-t-y$. For example, if $n_{p}=1$, we may take $z=\frac{1}{n}$ and $x_{i}\coloneq\sum_{j=0}^{i-1}z^{p^{j}}=\frac{i}{n}$ defines a sequence of period $p$, and $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{p^{i}}$ is a root of $t^{p}-t-y$. Because different choices of $n_{p}$ involve different choices for $z$ with $\textup{Tr}^{E}_{F}(z)=1$, this suggests that there are infinitely many (polynomial) formulas for the roots of polynomials $t^{p}-t-y\in E[t]$. If $E={\mathbb{F}}_{p^{n}}$, $F={\mathbb{F}}_{p}$ and $e=n_{p}$, then finding $z$ in ${\mathbb{F}}_{p^{e}}$ with $\textup{Tr}^{E}_{F}(z)=1$ is easy in practise: with probability $1-\frac{1}{p}$ a uniformly random element $\zeta\in{\mathbb{F}}_{p^{e}}$ has $\tau\coloneq\textup{Tr}^{E}_{F}(\zeta)\neq 0$, so $z=\tau^{-1}\zeta$ has $\textup{Tr}^{E}_{F}(z)=1$.
There are two versions of Hilbert’s Theorem 90: the additive form and the multiplicative form: we abbreviate these AH90 and MH90, respectively. Theorem 1.1 is different from usual version of AH90 in that the formula for $x$ is very slightly different. This difference leads to Theorem 1.2. We remark first that MH90 is so numbered because it appears as Satz 90 in [3] on p. 272, and Hilbert never actually proved AH90. The standard proof of AH90 is given in Lang [6, p. 215] and it follows Emil Artin’s ideas. In Theorem 1.1, $x$ equals $R(y,z)\coloneq\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\sigma^{j}z\sigma^{i}y$ where $z\in E$ satisfies $\textup{Tr}^{E}_{F}(z)=1$ and $y=\sigma x-x$ holds. By contrast, the standard proof of AH90 uses a different element, namely $x^{\prime}=R(z,y)$ (with $y,z$ swapped) and $y=x^{\prime}-\sigma x^{\prime}$ holds. The proofs of AH90 I have found, assume $\textup{Tr}^{E}_{F}(z)\neq 0$ (but scaling to get $\textup{Tr}^{E}_{F}(z)=1$ is easy) and show that
$x^{\prime}$ $=\frac{1}{\textup{Tr}^{E}_{F}(z)}\left(y\sigma(z)+(y+\sigma(y))\sigma^{2}(z)+\cdots+(y+\sigma(y)+\cdots+\sigma^{n-2}(y))\sigma^{n-1}(z)\right)$
$=\frac{1}{\textup{Tr}^{E}_{F}(z)}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\sigma^{j}y\sigma^{i}z=\frac{1}{\textup{Tr}^{E}_{F}(z)}R(z,y)$
satisfies $y=x^{\prime}-\sigma x^{\prime}$. The point here is that writing $x$ in the form $\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ for some $x_{i}\in E$, suggests studying the sequence $(x_{i})_{i\geqslant 0}$, whereas the form of $x^{\prime}$ does not.
We remark that the expression $x^{\prime}=R(z,y)$ is very natural in the context of Galois cohomology. Indeed, this form suggests itself via Emmy Noether’s equations, and these equations lead to analogous proofs of AH90 and MH90 as we explain. Emmy Noether generalized the method used to prove AH90 and MH90 by introducing 1-cocyles, to handle a noncyclic (finite) Galois group $\Gamma\coloneq\textup{Gal}(E/F)$, see [4, Theorems 4.29 and 4.30]. Her result [4, Theorem 4.29] says: given a function $y\colon\Gamma\to E\setminus\{0\}$ satisfying $y_{\beta\gamma}=y_{\beta}\beta(y_{\gamma})$ for $\beta,\gamma\in\Gamma$, then for some $z\in E$, $x=\sum_{\beta\in\Gamma}y_{\beta}\beta(z)$ satisfies $y_{\gamma}=x\gamma(x)^{-1}$ for all $\gamma\in\Gamma$. Her result [4, Theorem 4.30] says: given a function $y\colon\Gamma\to E$ satisfying $y_{\beta\gamma}=y_{\beta}+\beta(y_{\gamma})$ for $\beta,\gamma\in\Gamma$, then for some $z\in E$, $x=\textup{Tr}^{E}_{F}(z)^{-1}\sum_{\beta\in\Gamma}y_{\beta}\beta(z)$ satisfies $y_{\gamma}=x-\gamma(x)$ for all $\gamma\in\Gamma$. When $\Gamma=\langle\sigma\rangle$ is cyclic of order $n$, the 1-cocycle conditions imply that $y_{\sigma^{i}}=y(\sigma y)\cdots(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$, and similarly in the additive case $y_{\sigma^{i}}=y+(\sigma y)+\cdots+(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$. This gives rise to the multiplicative formula $x=\sum_{i=0}^{n-1}\left(\prod_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$ and when $\textup{Tr}^{E}_{F}(z)=1$, the additive formula $x=R(z,y)\coloneq\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$, respectively. Theorem 1.1 uses the related expression $R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$. The asymmetry between $R(y,z)$ and $R(z,y)$ is measured by the equation $R(y,z)+R(z,y)=-\sum_{i=0}^{n-1}\sigma^{i}(yz)$, see Lemma 3.1 below.
2 Background
This work fits into a very classical setting, namely generalizing Galois theory to characteristic $p$. Theorem 2.1 below follows from the the additive form (AH90) of Hilbert’s Theorem 90, it is proved in [4, Theorem 4.33] and [6, VIII, Theorem 11].
Theorem 2.1 (Artin-Schreier).
Let $E$ be field of prime characteristic $p$.
  1. (a)
    Let $E/F$ be a cyclic extension of degree $p$. Then $E=F(x)$ where $x^{p}-x\in F\setminus\{0\}$.
  2. (b)
    Given $y\in E$, the polynomial $t^{p}-t-y$ is either irreducible over $E$, or it has $p$ distinct roots $x+u$ where $x\in E$ and $u\in{\mathbb{F}}_{p}$.
The trace map $\textup{Tr}_{F}^{E}(x)=\sum_{\gamma\in\textup{Gal}(E/F)}\gamma(x)$ of a finite Galois extension $E/F$, is an $F$-linear map $E\to F$ which is nonzero by Artin’s Lemma [6, VIII, Theorem 7]. In certain cases, the trace map can be used to distinguish whether or not $t^{p}-t-y$ is irreducible.
Lemma 2.2.
Let $E/F$ be a cyclic Galois extension where $|E:F|=\textup{char}(F)=p$. Suppose that $\textup{Gal}(E/F)=\langle\sigma\rangle$ where $\sigma\colon E\to E$ is the $p$th power map $\sigma(x)=x^{p}$. For $y\in E$, the polynomial $t^{p}-t-y$ is reducible over $E$ if and only if $\textup{Tr}_{F}^{E}(y)=\sum_{i=0}^{p-1}y^{p^{i}}=0$. Furthermore, $E=F(z)$ for some $z\in E$ where $z^{p}-z^{p-1}+1=0$ and $\textup{Tr}^{E}_{F}(z)=1$.
Proof.
If $\textup{Tr}_{F}^{E}(y)=0$, then $y=\sigma(z)-z=z^{p}-z$ for some $z\in E$ by AH90. Hence $t^{p}-t-y$ is reducible over $E$. Conversely, suppose if $t^{p}-t-y$ is reducible over $E$, then $t^{p}-t-y$ splits over $E$ by Theorem 2.1(b). Let $z\in E$ be a root. Using $\sigma z=z^{p}$ shows
$0=\textup{Tr}^{E}_{F}(0)=\textup{Tr}^{E}_{F}(z^{p}-z-y)=\textup{Tr}^{E}_{F}(\sigma z)-\textup{Tr}^{E}_{F}(z)-\textup{Tr}^{E}_{F}(y)=-\textup{Tr}^{E}_{F}(y).$
Finally, set y=1y=-1. Then $\textup{Tr}_{F}^{E}(-1)=p(-1)=0$ and by the above argument $t^{p}-t+1=0$ is reducible over $E$. Hence some $\zeta\in E$ satisfies $t^{p}-t+1=0$ by Theorem 2.1(b). However, $\sigma\zeta=\zeta^{p}=\zeta-1$ implies $\sigma^{i}\zeta=\zeta-i$ so $\zeta$ has $p$ distinct conjugates in $E$ and $E=F(\zeta)$. Therefore $E=F(z)$ where $z=\zeta^{-1}$ is a root of the (reverse) polynomial $t^{p}(t^{-p}-t^{-1}+1)=t^{p}-t^{p-1}+1$. Since the conjugates $z,\sigma z,\dots,\sigma^{p-1}z$ are pairwise distinct, $t^{p}-t^{p-1}+1$ is irreducible over $F$. The coefficient of $t^{p-1}$ shows that $\textup{Tr}^{E}_{F}(z)=\sum_{i=0}^{p-1}\sigma^{i}z=1$. ∎
We note that only part of Theorem 2.1(b) generalizes. If ${\mathbb{F}}_{q}\leqslant F\leqslant E$, and $t^{q}-t-y$ has a root $x\in E$, then $t^{q}-t-y$ factors as $\prod_{a\in{\mathbb{F}}_{q}}(t-(x+a))$. However, if $t^{q}-t-y$ has no root in $E$, we can not conclude that $t^{q}-t-y$ is irreducible over $E$. For example, if $q=4$ then $t^{4}+t+1$ factors as $(t^{2}+t+\omega)(t^{2}+t+\omega^{2})$ over ${\mathbb{F}}_{4}={\mathbb{F}}_{2}(\omega)$.
A polynomial over a field of characteristic $p$ is solvable by radicals if its roots can be expressed as a finite combination of $\times,\div,\sqrt[r$]{\ \;} where $r$ is a prime different to $p$, and roots of Artin-Schreier polynomials [6, p. 217]. Note that every irreducible polynomial of degree $n$ over a finite field ${\mathbb{F}}_{q}$, is trivially solvable by radicals because each nonzero element of ${\mathbb{F}}_{q^{n}}$ is a primitive $m$th root of unity for some $m\mid(q^{n}-1)$. It makes sense, therefore, to allow non-radical expressions for the roots of polynomials over finite fields. One such example is $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$, which involves the ‘auxiliary’ element $z$.
3 Periodicity and Hilbert’s additive Theorem 90
Let $E/F$ be a Galois extension with group $\Gamma=\textup{Gal}(E/F)$. Thus $\textup{Tr}^{E}_{F}(\sigma x)=\textup{Tr}^{E}_{F}(x)$ for $\sigma\in\Gamma$. Hence if $y\in E$ has the form $\sigma x-x$ for some $x\in E$ and $\sigma\in\Gamma$, then $\textup{Tr}^{E}_{F}(y)=0$. If $\Gamma$ is finite and cyclic, then the converse holds by AH90. In this section $E/F$ is a cyclic extension of degree $n$, and $\textup{Gal}(E/F)=\langle\sigma\rangle$. Thus $\sigma^{n}=1$ and $\textup{Tr}_{F}^{E}(y)=\sum_{i=0}^{n-1}\sigma^{i}y$.
Proof of Theorem 1.1.
Suppose that $x$ has the form $\sum_{i=0}^{n-1}x_{i}\sigma^{i}y$ where $x_{0},\dots,x_{n-1}\in E$. We aim to solve for the coefficients $x_{i}$. Observe that
$\sigma x-x=-x_{0}y+\sum_{i=1}^{n-1}(\sigma x_{i-1}-x_{i})\sigma^{i}y+\sigma x_{n-1}\sigma^{n}y.$
Assume for the moment that there exists an element $z\in E$ satisfying
(1) $\sigma x_{i-1}-x_{i}=-z\qquad\textup{for $1\leqslant i<n$, and}\qquad\sigma x_{n-1}-x_{0}=1-z.$
Substituting these equations and using $\sigma^{n}y=y$ gives
$\sigma x-x=-x_{0}y-z\sum_{i=1}^{n-1}\sigma^{i}y+(1-z+x_{0})y=y-z\sum_{i=0}^{n-1}\sigma^{i}y=y$
as desired. Next, assume that $z$ and $x_{0}$ are given. The equations $\sigma x_{i-1}-x_{i}=-z$ show that $x_{i}=\sigma^{i}x_{0}+\sigma^{i-1}z+\cdots+\sigma z+z$ holds for $1\leqslant i<n$. Rewrite the $n$ equations (1) as $\sigma^{n-i+1}x_{i-1}-\sigma^{n-i}x_{i}=-\sigma^{n-i}z$ for $1\leqslant i<n$ and $\sigma x_{n-1}-x_{0}=1-z$. Adding these equations gives $0=\sigma^{n}x_{0}-x_{0}=1-\textup{Tr}_{F}^{E}(z)$. Hence $\textup{Tr}_{F}^{E}(z)=1$ must hold.
To see that these necessary equations are sufficient, choose $z\in E$ so that $\textup{Tr}_{F}^{E}(z)=1$ which is possible as the trace map $\textup{Tr}_{F}^{E}\colon E\to F$ is surjective. Next, define $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where the empty sum $x_{0}=\sum_{j=1}^{0}\sigma^{j}z$ is 0 by convention. Then the $n$ equations in (1) hold, and $\sigma x-x$ equals $y$, as claimed. ∎
Lemma 3.1.
Let $E/F$ be a cyclic Galois extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$. If $y,z\in E$ have traces $\textup{Tr}^{E}_{F}(y)=0$ and $\textup{Tr}^{E}_{F}(z)=1$, then $R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$ satisfies $R(y,z)=-\textup{Tr}^{E}_{F}(yz)-R(z,y)$, $\sigma R(y,z)-R(y,z)=y$ and $R(z,y)-\sigma R(z,y)=y$.
Proof.
The equations $\sum_{i=0}^{n-1}\sigma^{i}y=0$ and $\sum_{j=0}^{n-1}\sigma^{j}z=1$ imply
$R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y=\sum_{i=0}^{n-1}\left(1-\sum_{j=i}^{n-1}\sigma^{j}z\right)\sigma^{i}y=-\sum_{i=0}^{n-1}\left(\sum_{j=i}^{n-1}\sigma^{j}z\right)\sigma^{i}y.$
By swapping the order of summation, by swapping $i\leftrightarrow j$, and by using the formula $\textup{Tr}^{E}_{F}(zy)=\sum_{i=0}^{n-1}\sigma^{i}(zy)$, the previous display may be written
$R(y,z)=-\sum_{j=0}^{n-1}\sum_{i=0}^{j}\sigma^{j}z\sigma^{i}y=-\sum_{i=0}^{n-1}\sum_{j=0}^{i}\sigma^{i}z\sigma^{j}y=-\textup{Tr}^{E}_{F}(zy)-\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z.$
Therefore $R(y,z)=-\textup{Tr}^{E}_{F}(yz)-R(z,y)$. Theorem 1.2 proves that $\sigma R(y,z)-R(y,z)=y$ holds. Since $\sigma\textup{Tr}^{E}_{F}(yz)=\textup{Tr}^{E}_{F}(yz)$ it follows that $R(z,y)-\sigma R(z,y)=y$, as claimed. ∎
We prove in Theorem 1.2 that when $\textup{char}(F)=p>0$, the element $x$ in Theorem 1.1 has an intrinsic periodicity. This observation is used in Corollary 4.1 to give a polynomial formula for the roots of (reducible) Artin-Schreier polynomials.
Proof of Theorem 1.2.
(a) First, we show that the fixed subfield $K$ of $\sigma^{n_{p}}$ contains an element $z$ satisfying $\textup{Tr}^{E}_{F}(z)=1$. Let $n_{p^{\prime}}\coloneq n/n_{p}$ be the $p^{\prime}$-part of $n$. Since $n_{p^{\prime}}\neq 0$ in $F$ and the map $\textup{Tr}^{K}_{F}\colon K\to F$ is surjective, there exists an element $z\in K$ such that $\textup{Tr}^{K}_{F}(z)=(n_{p^{\prime}})^{-1}$. Therefore $\textup{Tr}^{E}_{K}(z)=|E:K|z=n_{p^{\prime}}z$ and hence $\textup{Tr}^{E}_{F}(z)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(z))=n_{p^{\prime}}\textup{Tr}^{K}_{F}(z)=1$. Next, suppose that $z\not\in K$ and $\textup{Tr}^{E}_{F}(z)=1$. Let $L=F(z)$. Then $1=\textup{Tr}^{L}_{F}(\textup{Tr}^{E}_{L}(z))=|E:L|\textup{Tr}^{L}_{F}(z)$. Therefore $p\nmid|E:L|$, so $n_{p}=|K:F|$ divides $e=|L:F|$ as claimed. In summary, the minimal polynomial over $F$ of an element $z$ with $\textup{Tr}^{E}_{F}(z)=1$ has degree a multiple of $n_{p}$, and degree precisely $n_{p}$ if $z$ lies in $K\setminus K_{0}$ where $K_{0}$ is the unique (maximal) subfield of $K$ with $F\leqslant K_{0}<K$ and $|K:K_{0}|=p$.
(b) Let $x_{i}=\sum_{j=0}^{i-1}\sigma^{j}z$ for $0\leqslant i<n$ where $\textup{Tr}^{E}_{F}(z)=1$. Let $L=F(z)$ and $e=|L:F|$. Hence $x_{e}=\textup{Tr}^{L}_{F}(z)\neq 0$ by part (b) and $\sigma^{j}z=z$ implies $e\mid j$. Suppose that $x_{k}=0$ where $0\leqslant k<pe$ and write $k=ie+j$ where $0\leqslant i<p$ and $0\leqslant j<e$. The equation $x_{\ell e+j}=\ell x_{e}+x_{j}$ shows that $x_{0}=x_{pe}=0$. Thus $x_{k}=ix_{e}+x_{j}=0$ and $x_{e}\in F$ gives
$x_{j+1}-z=\sigma x_{j}=\sigma(-ix_{e})=-ix_{e}=x_{j}.$
Whence $\sigma^{j}z=x_{j+1}-x_{j}=z$. This implies that $e\mid j$ and hence $j=0$. However, $0=x_{k}=x_{ie}=ix_{e}$ and $x_{e}\neq 0$ shows that $0=i=j=k$. Therefore $x_{\ell}\neq 0$ holds for $0<\ell<pe$, and the sequence $x_{0},x_{1},\dots$ is periodic, with period precisely $pe$. ∎
4 Formulas for the roots of Artin-Schreier polynomials
In this section we assume that $F$ has prime characteristic $p$. Set $q=p^{f}$. Two examples with $\textup{Gal}(E/F)=\langle\sigma\rangle$ cyclic of order $n$ and $\sigma(x)=x^{q}$ for $x\in E$, are the finite fields $E={\mathbb{F}}_{q^{n}}$, $F={\mathbb{F}}_{q}$, and the function fields $E={\mathbb{F}}_{q^{n}}(u)$, $F={\mathbb{F}}_{q}(u)$. We shall give a number of different applications of Theorem 1.2 depending on the value of $n_{p}$. Observe that an immediate corollary of Theorem 1.1 is a polynomial formula for a root of $t^{q}-t-y$.
Corollary 4.1.
Suppose that $\textup{char}(F)=p$, $q=p^{f}$, and $E/F$ is a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$ for $x\in E$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=0$ and $\textup{Tr}^{E}_{F}(z)=1$, then $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{q^{j}}y^{q^{i}}$ satisfies $x^{q}-x-y=0$.
When $n$ is coprime to $p$, i.e. $n_{p}=1$, the formula for $x$ in Corollary 4.1 simplifies.
Corollary 4.2.
Suppose that ${\mathbb{F}}_{q}\leqslant F$, $q=p^{f}$, and $E/F$ is a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\sigma(x)=x^{q}$ for $x\in E$. If $y\in E$, $\textup{Tr}^{E}_{F}(y)=0$ and $n_{p}=1$, then the factorization $t^{q}-t-y=\prod_{u\in{\mathbb{F}}_{q}}(t-x-u)$ holds where $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$.
Proof.
By Theorem 1.2(b), we may choose $z$ to lie in $F$ and so $1=\textup{Tr}^{E}_{F}(z)=nz$. Hence $z=\frac{1}{n}$ and $x=R(y,z)=\sum_{i=0}^{n-1}\frac{i}{n}y^{q^{i}}$ satisfies $x^{q}-x-y=0$ by Corollary 4.1. Therefore $x+u$ is a root of $t^{q}-t-y$ for each $u\in{\mathbb{F}}_{q}\leqslant F$, and the claimed factorization follows. ∎
Remark 4.3.
When $\textup{char}(F)=p\nmid n$, Corollary 4.2 implies that all the roots of $t^{q}-t-y$ lie in the ${\mathbb{F}}_{p}$-subspace ${\mathbb{F}}_{q}+\{\sum_{i=0}^{n-1}\lambda_{i}y^{q^{i}}\mid\lambda_{1},\dots,\lambda_{n-1}\in{\mathbb{F}}_{p}\}$ of ${\mathbb{F}}_{q^{n}}$.
When $n_{p}=p$, i.e. $p\mid n$ and $p^{2}\nmid n$, the element $z$ in Corollary 4.1 can be chosen to lie in the fixed field of $\sigma^{p}$, and to have a specific minimal polynomial.
Corollary 4.4.
Let $E/F$ be a cyclic field extension of degree $n$ where $n_{p}=\textup{char}(F)=p$. Suppose that $\textup{Gal}(E/F)=\langle\sigma\rangle$ where $\sigma\colon E\to E$ is the $p$th power map $\sigma(x)=x^{p}$. If $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}y^{p^{i}}=0$, and $\zeta$ satisfies $\zeta^{p}-\zeta^{p-1}+1=0$. Then $x^{p}-x-y=0$ where $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$ and $\zeta^{p^{p}-1}=1$.
Proof.
It follows from $n_{p}=p$, that $n/p$ is coprime to $p$. Let $K$ be the fixed field of $\sigma^{p}$. Then $|K:F|=p$ and $K=F(\zeta)$ where $\zeta$ satisfies $\zeta^{p}-\zeta^{p-1}+1=0$ and $\textup{Tr}^{K}_{F}(\zeta)=1$ by Lemma 2.2. Therefore $\textup{Tr}^{E}_{F}(\zeta)=n/p$ and $\textup{Tr}^{E}_{F}(z)=1$ where $z=(n/p)^{-1}\zeta$. By Theorem 1.1, $x=(n/p)^{-1}\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\zeta^{p^{j}}y^{p^{i}}$ satisfies $x^{p}-x-y=0$. Finally, $\sigma(\zeta)=\zeta^{p}$ implies that $\zeta=\sigma^{p}(\zeta)=\zeta^{p^{p}}$, so that $\zeta^{p^{p}-1}=1$. ∎
We shall now apply Corollary 4.1 in the case that $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$. In this case $E/F$ is cyclic and $\textup{Gal}(E/F)=\langle\sigma\rangle$ has order $n$ where $\sigma(x)=x^{p}$ is the $p$th power map.
Let $p,r$ be distinct primes. Denote by $\textup{ord}_{r}(p)$ the smallest positive integer $e$ such that $p^{e}\equiv 1\pmod{r}$ and note that $e\mid(r-1)$. The cyclotomic polynomial $\Phi_{r}(t)=\sum_{i=0}^{r-1}t^{i}$ of degree r1r-1 factors over ${\mathbb{F}}_{p}$ as a product of (r1)/e(r-1)/e irreducible polynomials each of degree $e=\textup{ord}_{r}(p)$ by [7, Theorem 2.47].
Theorem 1.2 shows that $z$ may be chosen to lie in a subfield of order $p^{f}$ where $f={n_{p}}$. Hence $z^{p^{f}-1}=1$. Lemma 4.7 proves that, in many cases, $z$ may be chosen to have prime order $r\neq p$. This choice of $z$, however, may increase the period of the sequence $(x_{i})_{i\geqslant 0}$.
Definition 4.5.
Call a non-zero polynomial $f(t)=\sum_{i=0}^{d}f_{i}t^{i}$ big if $d=0$ or $f_{d}f_{d-1}\neq 0$, and small if $f_{d}\neq 0$ and $f_{d-1}=0$. As a small polynomial times a small polynomial is small, and a big polynomial times a small polynomial is big, we have the following lemma.
Lemma 4.6.
If a big polynomial factors, then at least one of its factors is big.
Lemma 4.7.
Let $p,r$ be unequal primes. Let $e\coloneq\textup{ord}_{r}(p)$ divide $n$. Set $E={\mathbb{F}}_{p^{n}}$, $K={\mathbb{F}}_{p^{e}}$ and $F={\mathbb{F}}_{p}$. There exists a primitive $r$th root of unity $\zeta\in K$ for which $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. If $p\nmid\frac{n}{e}$ and $y\in E$ has $\textup{Tr}^{E}_{F}(y)=0$, then $x=(\frac{n\tau}{e})^{-1}\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\zeta^{p^{j}}\right)y^{p^{i}}$ satisfies $x^{p}-x-y=0$. Finally, the sequence $(\sum_{j=0}^{i-1}\zeta^{p^{j}})_{i\geqslant 0}$ is periodic with period $ep$ where $n_{p}\mid e$.
Proof.
First, $r$ divides $p^{d}-1$ if and only if $e\mid d$. As $e\mid n$, all primitive $r$th roots of unity lie in $K={\mathbb{F}}_{p^{e}}$, and $F\leqslant K\leqslant E$ holds. The cyclotomic polynomial $\Phi_{r}(t)=\sum_{i=0}^{r-1}t^{i}$ is big, and $\Phi_{r}(t)$ factors over $F$ as a product of irreducible polynomials, each of degree $e$, at least one of which is big by Lemma 4.6. Hence some $r$th root of unity $\zeta\in K$ has $\tau\coloneq\textup{Tr}^{K}_{F}(\zeta)\neq 0$. Thus $\textup{Tr}^{E}_{K}(\zeta)=\frac{n}{e}\zeta$ and $\textup{Tr}^{E}_{F}(\zeta)=\textup{Tr}^{K}_{F}(\textup{Tr}^{E}_{K}(\zeta))=\frac{n}{e}\textup{Tr}^{K}_{F}(\zeta)=\frac{n\tau}{e}\neq 0$. Therefore $z=(\frac{n\tau}{e})^{-1}\zeta\in K^{\times}$ satisfies $\textup{Tr}^{E}_{F}(z)=1$. The formula for $x=R(y,z)$ follows from Theorem 1.1 and the periodicity $ep$ and $n_{p}\mid e$ follows from Theorem 1.2(b). ∎
Lemma 4.8.
Let $p$ be a prime where $p\equiv 2\pmod{3}$. Let $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$ where $n$ is even and let $\omega\in E$ satisfy $\omega^{2}+\omega+1=0$. If $p\nmid(n/2)$, and $y\in E$ satisfies $\textup{Tr}^{E}_{F}(y)=0$, then the sum $x=(\frac{n}{2})^{-1}\sum_{i=0}^{n-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)y^{p^{i}}$ satisfies $x^{p}-x-y=0$ where $r_{i}=i-2\lfloor i/2\rfloor$. Finally, the sequence $x_{0},x_{1},\dots$ is periodic of period $2p$ where $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)$.
Proof.
Since $n$ is even and $p^{n}\equiv 1\pmod{3}$, $E$ contains a primitive cube root $\omega$ of 1. The condition $p\equiv 2\pmod{3}$ is equivalent to $e=\textup{ord}_{3}(p)=2$. Thus we may use Lemma 4.7 with $r=3$, $\zeta=\omega$ and $\tau=-1$. However, $\omega^{p^{k}}$ equals $\omega$ if $k$ is even, and $\omega^{2}$ if $k$ is odd. If $i=0,1,2,3,4,5,6$, then $s_{i}=\sum_{k=0}^{i-1}\omega^{p^{k}}$ equals 0, $\omega$, $\omega+\omega^{2}=-1$, $1+\omega$, 2-2, $2+\omega$, 3-3, etc. Therefore $s_{i}=-(i-r_{i})/2+r_{i}\omega$ where $r_{i}=i-2\lfloor i/2\rfloor\in\{0,1\}$. Consequently, the values of $x_{i}=(\frac{n}{2})^{-1}\left(\frac{i-r_{i}}{2}-r_{i}\omega\right)$ are periodic of period $2p$ by Theorem 1.2(b). ∎
The following result gives different quadratic ‘formulas’ over the finite field ${\mathbb{F}}_{2^{n}}$ that depend on $n_{2}$ and the choice of an element $z\in{\mathbb{F}}_{2^{n}}$ with absolute trace 1. The condition $n_{2}=2^{k}$ is equivalent to $n\equiv 2^{k}\pmod{2^{k+1}}$. The case $k=0$ is handled by Corollary 4.2.
Lemma 4.9.
Let $y\in{\mathbb{F}}_{2^{n}}$ where $\sum_{i=0}^{n-1}y^{2^{i}}=0$. Then $x=\sum_{i=0}^{n-1}x_{i}y^{2^{i}}$ is a root of $t^{2}+t+y$ where $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ and $z\in{\mathbb{F}}_{2^{n}}$ has absolute trace $\sum_{i=0}^{n-1}z^{2^{i}}=1$ and has degree $n_{2}$ and the sequence $(x_{i})$ has period $2n_{2}$. Table 1 lists choices for $z$ and the minimal polynomial $m_{z/{\mathbb{F}}_{2}}(t)$ of $z$ over ${\mathbb{F}}_{2}$, as a function of the $2$-part $n_{2}$ of $n$.
Table 1: Choices for $z$ and its minimal polynomial $m_{z/{\mathbb{F}}_{2}}(t)$ over ${\mathbb{F}}_{2}$ of degree $n_{2}$
$n_{2}$ 22 44 88 $16$ $32$
$z$ $\omega$ $\alpha$ $\beta$ $\gamma$ $\delta$
$m_{z/{\mathbb{F}}_{2}}(t)$ $t^{2}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{4}\kern-2.0pt+\kern-2.0ptt^{3}\kern-2.0pt+\kern-2.0pt1$ $t^{8}\kern-2.0pt+\kern-2.0ptt^{7}\kern-2.0pt+\kern-2.0ptt^{2}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{16}\kern-2.0pt+\kern-2.0ptt^{15}\kern-2.0pt+\kern-2.0ptt^{8}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$ $t^{32}\kern-2.0pt+\kern-2.0ptt^{31}\kern-2.0pt+\kern-2.0ptt^{3}\kern-2.0pt+\kern-2.0ptt\kern-2.0pt+\kern-2.0pt1$
Proof.
Set $E={\mathbb{F}}_{2^{n}}$, $F={\mathbb{F}}_{2}$ and $e=n_{2}$. The last line of Table 1 lists big irreducible polynomials over $F$ whose roots have order $2^{n_{2}}-1$, see Lemma 4.6 and Remark 4.10. The polynomials are big by Definition 4.5. The irreducibility over $F$, and the order of $z$, are easily checked using a computational algebra system. These together imply that $K\coloneq F(z)$ has $|K:F|=n_{2}$ and $\textup{Tr}^{K}_{F}(z)=1$. Hence $\textup{Tr}^{E}_{F}(z)=|E:K|=n_{2^{\prime}}=1$ as desired. By Theorem 1.1, $x=R(y,z)$ is a root of $t^{2}+t+y$. In order to use our formulas, the reader will need to know the values of $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$. Since $x_{0}=x_{2e}=0$ and $x_{i+2e}=x_{i}$, it suffices to list $x_{i}$ for $0\leqslant i<2e$. However, since $x_{e}=\textup{Tr}_{F}^{K}z=1$ and $x_{i+e}=1+x_{i}$, it suffices to list $x_{i}$ for $0\leqslant i<e$.
(a) Suppose that $n_{2}=1$. Set $q=p=2$ in Corollary 4.2. Then $x^{2}+x+y=0$ where $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{2^{i}}=\sum_{j=0}^{(n-1)/2}y^{2^{2j+1}}$. Here $x_{i}\in\{0,1\}$ satisfies $x_{i}\equiv i\pmod{2}$.
(b) Suppose that $n_{2}=2$. Set $p=2$ and $z=\omega$ into Lemma 4.8. Since $n/2$ is odd the value of $x_{i}=(\frac{n}{2})^{-1}\left(\lfloor\frac{i}{2}\rfloor-r_{i}\omega\right)$ is $0,\omega,1,\omega^{2}$ when $i\equiv 0,1,2,3\pmod{4}$; recall that $r_{i}=i-2\lfloor\frac{i}{2}\rfloor\in\{0,1\}$. Equivalent formulas for $x$ in cases (a) and (b) were given previously by Chen [1, Theorems 1, 2] and by Lahtonen [5].
(c) Suppose that $n_{2}=4$. Then $z=\alpha\in{\mathbb{F}}_{16}$ has order $2^{4}-1$ and $\textup{Tr}^{K}_{F}(z)=1$. Also the values of $x_{i}=\sum_{j=0}^{i-1}z^{2^{i}}$ are $0,\alpha,\alpha^{13},\alpha^{6},1,\alpha^{12},\alpha^{7},\alpha^{8}$ for $i\equiv 0,1,\dots,7\pmod{8}$.
(d) Suppose that $n_{2}=8$. Then $z=\beta\in{\mathbb{F}}_{256}$ has order $2^{8}-1$ and $z$ satisfies $\textup{Tr}^{K}_{F}(z)=1$. Furthermore, the values of $x_{i}=\sum_{j=0}^{i-1}z^{2^{i}}$ are $0,\beta,\beta^{100},\beta^{189},\beta^{29},\beta^{60},\beta^{154},\beta^{177}$ for $i\equiv 0,1,\dots,7\pmod{16}$. Also $x_{i+8}=1+x_{i}$ for $i\equiv 0,1,\dots,7\pmod{16}$.
(e) Suppose that $n_{2}=16$. Then $z=\gamma\in{\mathbb{F}}_{2^{16}}$ has order $2^{16}-1$ and $\textup{Tr}^{K}_{F}(z)=1$. In addition, $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ equals $0,\gamma,\gamma^{64409},\gamma^{48754},\gamma^{27742},\gamma^{48469},\gamma^{1146},\gamma^{22404},\gamma^{64313},\gamma^{47682},$ $\gamma^{63219},\gamma^{45929},\gamma^{55680},\gamma^{46875},\gamma^{7495},\gamma^{32204}$ and $x_{i+16}=1+x_{i}$ for $i\equiv 0,1,\dots,15\pmod{32}$.
(f) Suppose that $n_{2}=32$. Then $z=\delta\in{\mathbb{F}}_{2^{32}}$ has order $2^{32}-1$ and $\textup{Tr}^{K}_{F}(z)=1$. Writing $x_{i}=\sum_{j=0}^{i-1}z^{2^{j}}$ as a power of $\delta$ for $i\equiv 0,1,\dots,15\pmod{32}$, involves solving the discrete logarithm problem. This was done in the previous cases as writing a power of a primitive element was shorter than writing polynomials in the primitive element. ∎
Remark 4.10.
It is an open question whether big primitive polynomials of degree $e=n_{2}=2^{k}$ in the last line of Table 1 exist for arbitrarily large choices of $k$. This turns out to be true if $2^{e}-1$ is square-free because by Lemma 4.11 the cyclotomic polynomial $\Phi_{2^{e}-1}(t)$ is big. This follows because $\Phi_{2^{e}-1}(t)$ is a tensor product [2] $\Phi_{p_{1}}(t)\otimes\cdots\otimes\Phi_{p_{r}}(t)$ of big polynomials where $2^{e}-1=p_{1}\cdots p_{r}$ and the $p_{i}$ are distinct primes. However, it is not known whether $2^{2^{k}}-1=\prod_{i=0}^{k-1}(2^{2^{i}}+1)$ is square-free for all $k\geqslant 1$. (The Fermat numbers $2^{2^{i}}+1$ are pairwise coprime, but it not known whether Fermat numbers are square-free.) The requirement that the polynomial be primitive (i.e. a root $z$ has order $2^{e}-1$) is not necessary, we assumed primitivity only so that the non-zero elements $x_{i}$ could be written as a power of $z$. What surprised the author is that a root of the Schreier polynomial $t^{2}+t+y$ in ${\mathbb{F}}_{2^{n}}$ can always be described using only ${\rm O}(e^{2})$ bits of information.
Lemma 4.11.
The tensor product of big nonzero polynomials is big, see Definition 4.5.
Proof.
Let $a(t)$ and $b(t)$ be big nonzero polynomials of respective degrees $m$ and $n$. The result is clear if $mn=0$ by [2, p. 307]. Suppose that $mn>0$ and $a(t)=\prod_{i=1}^{m}(t-\alpha_{i})$ and $b(t)=\prod_{j=1}^{n}(t-\beta_{j})$ are monic with roots $\alpha_{i},\beta_{j}$ in an extension field. By Definition 4.5, $a(t)=t^{m}+a_{m-1}t^{m-1}+\cdots$ and $b(t)=t^{n}+b_{n-1}t^{n-1}+\cdots$ where $a_{m-1}=-\sum_{i=1}^{m}\alpha_{i}\neq 0$ and $b_{n-1}=-\sum_{j=1}^{n}\beta_{j}\neq 0$. However, $(a\otimes b)(t)=\prod_{i,j}(t-\alpha_{i}\beta_{j})$ by [2, p. 307], and hence $(a\otimes b)(t)=t^{mn}-(\sum_{i,j}\alpha_{i}\beta_{j})t^{mn-1}+\cdots$. Therefore $(a\otimes b)(t)$ is big since $\sum_{i,j}\alpha_{i}\beta_{j}=a_{m-1}b_{n-1}\neq 0$. More generally, if $a(t)$ and $b(t)$ have leading terms $a_{m}$ and $b_{n}$ respectively, then $(a\otimes b)(t)$ has leading term $a_{m}^{n}b_{n}^{m}\neq 0$ by [2, p. 307], and hence is big. ∎